### v1

# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for Foundation Models

Abstract.
The burgeoning field of foundation models necessitates advanced data processing mechanisms capable of harnessing vast valuable data with varied types utilized by these models. Nevertheless, the current landscape presents unique challenges that traditional data processing frameworks cannot handle effectively, especially with multimodal intricacies. In response, we present Data-Juicer 2.0, a new system offering fruitful data processing capabilities backed by over a hundred operators spanning various modalities like text, image, audio, and video. With seamless compatibility and dedicated optimization to popular dataset hubs like Hugging Face and computing engines like Ray, Data-Juicer 2.0 enhances its predecessor in both usability, efficiency, and programmability. It features an easily accessible user interface layer that supports decoupled Python interactions, RESTful APIs, and conversational commands. Alongside this, it contains a core runtime layer optimized for adaptive execution and management across different dataset scales, processing demands, and computational environments, while shielding unnecessary system details. Extensive empirical evaluations demonstrate Data-Juicer 2.0’s remarkable performance and scalability, highlighting its capability to efficiently process tens of billions of data samples with tens of thousands of CPU cores. The system is publicly available, actively maintained, and broadly adopted in diverse research endeavors, practical applications, and real-world products such as Alibaba Cloud PAI.
11footnotetext: Co-first authors.22footnotetext: Equal contribution.33footnotetext: Corresponding authors, email addresses: {yaliang.li, bolin.ding}@alibaba-inc.com
1. Introduction
Figure 1. The overview of Data-Juicer 2.0.
Data Processing Challenges for Foundation Models. The advent of foundation models capable of processing textual and multimodal information, exemplified by the GPT-series (OpenAI, 2024), has transformed artificial intelligence by harnessing expansive datasets and substantial computational resources (Wang et al., 2023b; Zhao et al., 2023). The efficacy of these models depends on meticulous data collection, cleaning, annotation, and synthesis (Wang et al., 2023b; Qin et al., 2024b; Bai et al., 2024b), essential for optimizing performance, scalability, and robustness across complex applications.
However, current data processing frameworks, whether for big data (Zaharia et al., 2010) or text-only foundation models (Computer, 2023; Soldaini, Luca and Lo, Kyle and Kinney, Rodney and Naik, Aakanksha and Ravichander, Abhilasha and Bhagia, Akshita and Groeneveld, Dirk and Schwenk, Dustin and Magnusson, Ian and Chandu, Khyathi, 2023; Jennings et al., [n.d.]), are inadequate for handling the complexities of multimodal datasets, due to unique challenges as the field rapidly evolves:
Practitioners lack open-source frameworks specifically designed to manage the diverse data types and semantics integral to foundation models. The processing of these datasets extends beyond standard operations to include specialized tasks such as semantic understanding and cross-modal alignment, which are crucial for optimizing the performance and utility of foundation models (Li et al., 2023b; Tong et al., 2024). For example, the inaugural version, Data-Juicer 1.0  [^1] In this paper, references to Data-Juicer 1.0 pertain to the SIGMOD publication (Chen et al., 2024a) and the open-source codes versioned with the v0.1.2 tag., provided a unified data processing framework for language models with approximately 50 text-only operators (OPs) (Chen et al., 2024a). As user demands and applications evolved rapidly, its complexity increased correspondingly. The need for heterogeneous data processing spanning images, audio, and video underscores the necessity for Data-Juicer 2.0 with enhanced usability, calling for an updated version that can provide advanced coordination mechanisms across various modules, OPs, resources, and models utilized in data processing.
Furthermore, foundation models set specific efficiency demands on data pipelines, emphasizing simple yet computationally intensive operations that traditional big data engines struggle to perform (Team, 2024a; Jennings et al., [n.d.]; Lhoest et al., 2021). In Data-Juicer 1.0, the tightly coupled executors and OP functions hindered stability and efficiency optimization when handling complex, large-scale datasets. In scenarios where a defect is discovered late in handling a massive 10TB dataset at the 9.9TB mark, recovery and cost reduction present formidable challenges.
Moreover, there is a shift towards user interfaces that cater to Python-based ecosystems, preferred by foundation model practitioners (Paszke et al., 2019; Abadi et al., 2016; Rasley et al., 2020), raising additional potential efficiency issues with Python. Addressing new practical needs, Data-Juicer 1.0 showed shortcomings in programmability, such as fragmented development workflows across diverse underlying frameworks like Hugging Face Dataset (Lhoest et al., 2021) and Ray Dataset (Moritz et al., 2018). An ideal solution would enable automatic adaptation to different data modalities, processing requirements, and software-hardware environments, ensuring a smooth, user-friendly development and data processing experience.
Overview of the New Architecture. To address these challenges and respond to emerging needs, we present Data-Juicer 2.0 as illustrated in Fig. 1, emphasizing new functionality and adaptability across hierarchical layers:
Data Processing Capability Layer. Data-Juicer 2.0 provides a broader and more advanced array of capabilities that encompass hundreds of OPs, responding to evolving data processing needs across a wide array of foundation model applications. While Data-Juicer 1.0 focused primarily on data cleansing and lightweight algorithmic implementations, Data-Juicer 2.0 expands the scope beyond the text domain to incorporate image, audio, and video modalities with more functionalities. It offers over 80 new OPs for data analysis, cleaning, synthesis, augmentation, and deduplication, utilizing semantic-aware auxiliary models like GPT and advanced algorithms such as load-balanced distributed union-find. It also includes more compositional OP types to enhance expressive capabilities. Additionally, integrating seamlessly with dataset hubs from Hugging Face (Team, 2024f) and ModelScope (Team, 2024g), it is compatible with massive existing multimodal datasets.
User Interface Layer. Offering a versatile suite of interfaces, Data-Juicer 2.0 caters to various user requirements across multiple operational and abstraction levels. Advanced developers seeking extensive customization will benefit from systematically structured and uniformly designed low-level Python APIs. For simplified service deployment and application integration, lightweight RESTful APIs are available, complemented by a web-based graphical interface that facilitates intuitive and rapid data workflow construction via drag-and-drop features. Moreover, leveraging recent advancements in large language models, natural language processing capabilities are introduced to allow users to issue data processing commands in conversational formats, enhancing accessibility and ease of use.
Core Runtime Layer. At Data-Juicer 2.0’s core is a multi-tiered core runtime optimized for flexible and adaptive processing. It is designed and implemented into four main parts:
(1) A unified Data-Juicer-Dataset (DJ-Dataset) class provides user-friendly programming abstractions and standard capability interfaces across diverse dataset classes, including Hugging Face Dataset, Ray Dataset, and Alibaba Cloud’s MaxCompute DataFrame (Cloud, 2024c), suitable for both standalone and cloud-scale environments;
(2) Decoupled Data-Juicer-Operators (DJ-Operators) classes, which deliver foundational primitives with flexible invocation, automatic optimization, and enhanced capabilities for inspection, integration, and testing. Various design principles and patterns are incorporated (Gamma, 1995), such as Single Responsibility, Strategy, Decorator, Abstract Factory, and Template Method;
(3) Control-panel components orchestrate and manage data processing jobs. These components, like the executor, integrate DJ-Dataset, DJ-Operators, monitors, and other modules seamlessly, forming templates and workflows with proven effectiveness. We extend them beyond dataset-specific development to support data-model co-development (Chen et al., 2024c), utilizing cost-effective model training and evaluation signals via an experimental sandbox;
(4) Adapter features are applied to both DJ-Dataset and DJ-Operators to enable automatic adaptation of data processing tasks to variations in computational requirements (e.g., different OP orchestrations and workloads), software environments (various data engines), and hardware configurations (varying memory, CPU cores, and GPUs). Internally, the Dataset Adapter includes features like engine-agnostic compatibility, streaming I/O, and fault tolerance, while the Operator Adapter facilitates automatic workload-aware OP reordering, OP-wise parallelism, and insight mining.
Contributions. Our contributions are summarized below:
Enhanced System and Adaptive Techniques. Incorporating feedback from numerous users and cutting-edge applications built upon its predecessor, Data-Juicer 2.0 emerges as a new open-source system with enhanced multimodal processing capabilities. It emphasizes dedicated multi-layered adaptability and system optimization, integrating cloud-scale practices and diverse computation engines to dynamically and efficiently meet data processing demands.
Extensive Evaluation and Usages. We conduct thorough experimental analyses to assess the system’s capabilities under varied workloads, such as Filter-Heavy, Mapper-Heavy, Model-Heavy, Multimodal, and Fuzzy-Deduplication scenarios, using datasets ranging from millions to tens of billions of entries. Our findings provide detailed insights and illuminate performance trade-offs across different use cases and resources, involving Ray and Alibaba MaxCompute with up to 100 nodes and 12,800 cores.
Community and Applications. We have open-sourced the new Data-Juicer 2.0 system at https://github.com/modelscope/data-juicer, fostering sustained community maintenance and engagement through practical events, such as tutorials (Chen et al., 2024b), surveys (Qin et al., 2024b), and data competitions (Team, 2024e). Data-Juicer 2.0 also facilitates many foundation model researches (Ge et al., 2024; Qin et al., 2024a; Bai et al., 2024a; Jiao et al., 2024) and serves as the operator base for multiple Alibaba Cloud products such as PAI Designer (Cloud, 2024d), indirectly benefiting hundreds of internal and external customers across various real-world AI businesses. One of them has been running stably for over five months, processing data at a scale exceeding terabytes.
Organization. The remainder of this paper is structured as follows. We first elaborate on the identified challenges in data processing for foundation models (Sec. 2.1), and outline the corresponding design objectives for Data-Juicer 2.0 (Sec. 2.2). In Sec. 3, we examine the operational scope and practical utility of Data-Juicer 2.0. In Sec. 4, we present various user interfaces, ranging from low-level APIs to language-based interactions. In Sec. 5, we discuss the design of DJ-Dataset (Sec. 5.1), DJ-Operators (Sec. 5.2), control components (Sec. 5.3), and adapter features for datasets (Sec. 5.4) and OPs (Sec. 5.5). Finally, Sec. 6 delivers comprehensive evaluations across diverse scales, and we conclude in Sec. 7.
2. Background
2.1. Data Processing for Foundation Models
Foundation models, especially those beyond text, demands innovative data processing paradigms to address the complexity and diversity of large-scale training datasets (Qin et al., 2024b; Zhao et al., 2023; Wang et al., 2023b; Bai et al., 2024b). Despite existing systems for big data or text-only models (Zaharia et al., 2010; Computer, 2023; Soldaini, Luca and Lo, Kyle and Kinney, Rodney and Naik, Aakanksha and Ravichander, Abhilasha and Bhagia, Akshita and Groeneveld, Dirk and Schwenk, Dustin and Magnusson, Ian and Chandu, Khyathi, 2023; Jennings et al., [n.d.]), there is a notable lack of open-source frameworks customized for foundation models due to unique design and implementation challenges:
Functionality Challenges from Data Types and Semantics. Data processing frameworks for foundation models primarily target semi-structured textual and multimodal datasets (e.g., JSONL, Parquet, MP4), often disregarding the strongly structured data types prominent in previous big data systems (e.g., database tables) (Lhoest et al., 2021; Jennings et al., [n.d.]). Significantly, the performance of foundation models depends on the quality, diversity, utility, and safety of data, necessitating a profound semantic understanding and cross-modal alignment to transform raw data into high-quality inputs, thereby enhancing model efficacy (Liu et al., 2023; Li et al., 2023b; Tong et al., 2024).
Efficiency Challenges from Data Operations and Pipelines. Conventional engines support a broad range of SQL queries for complex data operations (Zaharia et al., 2010; Team, 2024a). However, in the context of foundation models, the requirement for traditional grouping and aggregation operations is relatively minimal (Bai et al., 2024b). Instead, emphasis is placed on simpler operations such as Mapper and Filter, which rely heavily on deep learning models via local inference (Schuhmann et al., 2021; Byeon et al., 2022; Wang et al., 2023a) or API calls (Liu et al., 2023; Mei et al., 2024; Maaz et al., 2023). Furthermore, individual records in these datasets can be exceedingly large, surpassing the capabilities of traditional big data engines (Naeem et al., 2022; Ikegwu et al., 2022). When combining multiple data operations into pipelines, it is crucial to maintain efficient data and transformation flows throughout the data lifecycle to prevent bottlenecks stemming from computational overhead, latency, and data integrity.
Usage Challenges from User Interfaces. Traditional big data systems typically use programming interfaces based on Java and Scala, supplemented by SQL. However, practitioners focusing on foundation models predominantly operate within the Python ecosystem, supporting both upstream and downstream repositories and workflows (Paszke et al., 2019; Abadi et al., 2016; Rasley et al., 2020). This shift necessitates designing user interfaces that specifically cater to the new programming needs of foundation model practitioners.
2.2. System Design Goals in Data-Juicer 2.0
Built upon Data-Juicer 1.0, we aim to address challenges identified in Sec. 2.1 with the focus on the following new objectives:
Comprehensive Multimodal Processing Abilities. The first basic goal is to support extensive multimodal processing capabilities across diverse domains, such as AI-generated content, spatial intelligence, education, and entertainment. This involves meticulous processes such as collection, cleansing, annotation, and synthesis, which integrate perceptual and cognitive information across modalities such as video, image, text, and audio (Brooks et al., 2024; Podell et al., 2023; Achiam et al., 2023; Copet et al., 2023).
Efficient Data Processing Operators and Data Flow. Comprehensive semantic operations are expected through advanced procedures that may be high-cost, such as deep learning model invocations, fuzzy deduplication, and clustering. For basic, high-frequency operations, we need to employ dedicated code optimization and speedup techniques like GPU computation or Cython (Behnel et al., 2011). Additionally, optimizing the loading, parsing, and storage of multimodal data, while minimizing overhead associated with linking, read/write operations, and network transmissions, is crucial for streamlining processing and enhancing efficiency.
Easy-to-Use Interface. To tackle usage challenges, Data-Juicer 2.0 pursues an intuitive, engine-agnostic abstraction of computational primitives, supporting Python, layered APIs, graphical user interfaces, and even natural language interactions. This design shields users from the inherent complexities of various computational engine implementations and manages the cognitive and usage costs to remain consistent, even as functionalities expand.
Adaptive and Scalable Processing. Last but not least, Data-Juicer 2.0 is designed to dynamically adapt to diverse computational environments and workloads, optimizing performance across frameworks including Hugging Face Dataset, Ray Dataset, and Alibaba Cloud MaxCompute, while considering their respective use cases. For example, OPs based on Hugging Face Dataset align well with large model community compatibility but require enhancements for scalable and distributed processing. Ray-based OPs offer excellent Python compatibility with low interaction overhead, yet they may face single-node bottlenecks for large-scale clusters involving data shuffles, such as those required by join operation. OPs based on Spark and MaxCompute necessitate careful resource management and schema considerations due to serialization between Python and the backend engines (e.g., JVM), rendering them less suitable for high-frequency computations.
In the subsequent sections, we will explore the design specifics and implementation of Data-Juicer 2.0 to illustrate how these objectives are achieved.
3. Processing Capabilities Beyond Text
3.1. Overview of New Operators
Data-Juicer 2.0 extends its predecessor, Data-Juicer 1.0—which included approximately 50 text-only OPs—by incorporating over 80 new OPs crafted for multimodal data processing with more functionalities, encompassing modalities such as image, video, and audio. A detailed list of these OPs is available in Table LABEL:tab:mm_ops in the Appendix. For clarity and ease of reference, the OPs are categorized from the following aspects and visually represented in Fig. 2.
  • Modality Types: The majority of new OPs are concentrated on video/image/text-only processing, with approximately 20 OPs dedicated to cross-modal data processing.
  • Operator Types: About 90% are Mappers and Filters, reflecting the user needs in the foundation model domain. New variants in other types, such as Formatters and Deduplicators remain stable due to the broad applicability of already supported formats (e.g., JSONL, Parquet, MP4) and classical algorithms like MinHash (Broder, 1997).
  • Function Types: Data cleaning and analysis operations constitute about half of the new OP suite. Notably, Data-Juicer 2.0 introduces about 40 OPs for data synthesis for cross-modal and post-tuning scenarios, enabling any-to-any generation among 4 supported modalities, and enhancing textual instruction and response based on varied information and needs. Furthermore, new privacy protection OPs and several wrapper OPs for established tools, such as FFmpeg (Developers, [n.d.]), have been integrated, enabling users’ convenient invocation of traditional professional commands.
  • Implementation Types: Model-based OPs dominate, since semantic-aware processing often requires advanced models, such as large models for captioning and visual analysis via SDXL (Podell et al., 2023), GPT (Achiam et al., 2023), and Qwen (Team, 2024c).
(a) Modality type.
(b) Operator type.
(c) Function type.
(d) Implementation type.
Figure 2. Distribution of new OPs across various dimensions.
3.2. Showcase of Typical New Operators
The new OPs include some unique contributions to Data-Juicer 2.0, and others partly inspired by state-of-the-art data processing methodologies for foundation models (Schuhmann et al., 2021; Zhu et al., 2023; Gadre et al., 2024). Below are three representative examples that showcase the diverse computational operations and requirements of these OPs.
phrase_grounding_recall_filter: This OP, newly developed by Data-Juicer 2.0, assesses alignment and consistency between images and textual descriptions. As illustrated in Fig. 3a, it identifies noun phrases in the text that refer to key entities. Subsequently, an open-vocabulary object detection model, such as Owl-ViT (Minderer et al., 2022), attempts to detect corresponding entities within the image. The OP then calculates the recall of detected phrases to evaluate consistency, where a higher recall indicates better image-text coherence.
video_motion_score_filter: This OP quantifies video dynamics. As shown in Fig. 3b, it samples multiple frames at a specified frames-per-second (FPS) rate and computes optical flows. The average magnitude of these flows determines the motion score, with higher scores indicating more dynamic content. To accommodate various computational resources, both a GPU-based RAFT version (Teed and Deng, 2020) and a CPU-based OpenCV version (Bradski, 2000) are available.
video_captioning_from_summarizer_mapper: This OP generates new video captions by combining tagging and captioning capabilities from various OPs. As illustrated in Fig. 3c, it uses six OPs from different angles: two for tagging and captioning the audio stream, two for static visual frame analysis, and two for dynamic video stream and information integration. A summarizer finally incorporates the three captions and two tag sets into a new caption. This composition yields more accurate and comprehensive captions by considering multi-dimensional and multi-perspective content.
(a) Example of phrase_grounding_recall_filter.
(b) Example of video_motion_score_filter.
(c) Example of video_captioning_from_summarizer_mapper.
Figure 3. Illustrative examples of new operators.
4. User Interfaces
4.1. Low-Level APIs
The framework’s core capabilities are exposed through Python-based programmatic interfaces, providing object-oriented logical encapsulation for both the fundamental DJ-Dataset, DJ-Operators and other runtime modules. This design delivers developers maximum flexibility and customizability. Processing workflows can be automatically chained by passing a series of OP instances to a loaded DJ-Dataset object (e.g., data.process([op1, op2])), enabling various operations to be performed on the dataset in a single pass. Additionally, the framework supports applying an instantiated DJ-Operators to a target dataset (e.g., op.run(data)), enhancing the reusability of OP instances. More details about the DJ-Dataset and DJ-Operators will be introduced later in Sec. 5. This dual approach—chained processing and individual OP application—optimizes both efficiency and modularity in data processing tasks, while leveraging the inherent advantages of Python’s ecosystem.
4.2. RESTful APIs
Utilizing standard Python type hints, we provide one-click generation of high-performance web APIs, capable of automatically discovering, registering, and adapting OP classes and related tools. Users can rapidly initiate a web server supporting the Asynchronous Server Gateway Interface by simply executing a service script, eliminating the need for manual code writing. The asynchronous concurrency mechanism enables options such as lightweight background tasks and mitigates potential bottlenecks for endpoints that may experience prolonged network I/O blocking. Each OP is accessible via POST requests, typically executing the OP’s run() method as the endpoint. The target dataset path is passed through query parameters, with additional configurable OP parameters transmitted via JSON payload. Upon completion, the path to the processed dataset is returned. This invocation through Web APIs allows for the centralized host with distributed access, reducing deployment complexity. It also facilitates the separation of application logic from execution, potentially fostering the development and release of more applications built upon Data-Juicer. Importantly, the extensive customization parameters available in the programming API can be seamlessly passed through the Web API, maintaining full functionality without compromising usability, and facilitating serviceful calling by higher-level interfaces as introduced later.
4.3. Web & CMD Tools
Figure 4. An example of config file for data recipe.
Thanks to the low-level and RESTful APIs, we further establish service-level capabilities across various scenarios in Data-Juicer 2.0, encapsulating scenario-specific processing solutions into built-in data recipes, described as YAML configuration files for end-to-end workflows (an example for video-data-synthesis is shown in Fig. 4). Centered around these data recipes, we provide user-friendly, highly encapsulated command-line interfaces as illustrated in Listing 4.3, supporting automated lightweight installation and recipe-level invocation. These tools automatically gather and install the requisite of OPs within the recipe, perform automated OP fusion, and maximize adaptive utilization of computing resources while minimizing user cognitive load and operational cost.
{listing}
[!ht] {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]bash dj-install –config recipe.yaml # install dependencies only for this recipe dj-analyze –auto –dataset_path xx.jsonl # default OPs, or via specific recipe dj-process –config recipe.yaml # for both standalone or distributed Representative tools available in Data-Juicer 2.0.
Utilizing the aforementioned APIs and tools, Data-Juicer’s capabilities are seamlessly integrated into Alibaba Cloud’s visual modeling product, PAI-Designer (Cloud, 2024d). As shown in Fig. 1, DJ-Operators are encapsulated as visual components within Designer, presented in a directory tree structure for easy access. Users can manage data by intuitive drag-and-drop manners, and import built-in or customized recipes in YAML format using graphical canvas interface. With Designer as entry point, users can activate MaxCompute cluster nodes (Cloud, 2024c) or containerized deep compute clusters (PAI-DLC) (Cloud, 2024b) directly for efficient cloud-native distributed data processing, leveraging Alibaba Cloud’s data storage solutions (e.g., Object Storage OSS (Cloud, 2024f), Cloud Parallel File System CPFS (Cloud, 2024a), and NAS (Cloud, 2024e)) along with automated performance monitoring and optimization.
4.4. Natural Language Interaction
User:: Filter out too short text samples. The dataset path is input_data.jsonl Assistant:: <thought>Filter out text samples that are too short based on the provided dataset path.</thought>
<function>run_text_length_filter</function>
<dataset_path>input_data.jsonl</dataset_path>
[…]
[ARGUMENTS] dataset_path: input_data.jsonl
[STATUS] SUCCESS
[RESULT] Filtered dataset path: output.jsonl
Assistant:: <thought>The filtering has been successfully executed, and the filtered dataset is now at the specified path.</thought>
<function>finish</function>
<response>Your request to filter out too short text samples has been completed. The processed dataset is available at output.jsonl.</response>
Assistant:: Your request to filter out too short text samples has been completed. The processed dataset is available at output.jsonl.
Querying and understanding over 100 diverse OPs have become challenging. Thankfully, advances in foundation models have transformed human-computer interactions. Research shows the synergy of reasoning and acting (ReAct) (Yao et al., 2022) in language models enhances task-solving. Using RESTful API services, we developed model-driven agents based on ReAct, enabling intuitive data processing through natural language. We adopted AgentScope (Gao et al., 2024), a multi-agent platform, for low-code integration, using prompt-optimized ReAct agents to align OP functionalities with our RESTful APIs. Built-in function preprocessing, response parsing, and fault-tolerance streamline development. Data-Juicer 2.0 thus allows flexible instruction input, interpreting user intents to execute appropriate data processing functions.
An example is shown above for interactive text length filtering using this agent, where users describe tasks with simple and vague language. Specifically, the user requests “filter out too short text samples from input_data.jsonl”. The agent’s ReAct prompt helps analyze and reason user intent, signaled by the <thought> tag. The <function> tag notes the chosen function, while <dataset_path> represents parameter mapping. Initially, a statistical analysis of text lengths is visualized, and the agent asks for filtering thresholds. After user input, filtering runs with parameters via RESTful API. Upon completion, the agent confirms the processed dataset path as output.jsonl, marks the task finished, and informs the user. Gray dialogs reveal the agent’s reasoning and actions, providing insights into the workflow. This example illustrates step-by-step user understanding, function selection, parameter population, and task execution, emphasizing transparency and automation of Data-Juicer 2.0 with agentic processing.
5. Data-Juicer Core Runtime
5.1. Data-Juicer-Dataset
5.1.1. Programming Interfaces with Extensive Compatibility
The core runtime of Data-Juicer 2.0 is implemented in Python, featuring a flexible data processing pipeline characterized by two execution modes: a standalone mode for convenient single-node execution, based on the Hugging Face Dataset (Lhoest et al., 2021), and a distributed mode that offers scalability across multiple nodes, leveraging the Ray Data (Moritz et al., 2018) and MaxFrame-DataFrame (Cloud, 2024c). These frameworks offer diverse computational engines, each with distinct capabilities and suitable use cases, but they also come with heterogeneous programming interfaces and complex implementation details.
In Data-Juicer 2.0, we exploit the strengths of these diverse computational engines and their associated programming classes by abstracting a top-level DJ-Dataset class and devising a standardized data representation schema (detailed in the next subsection). The primary design principle employed is the Facade Pattern. The abstracted class provides comprehensive and unified development interfaces, facilitating seamless use across standalone, distributed, and cloud cluster environments, meanwhile optimized to support extensive multimodal data processing.
{listing}
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]python # Differernt input sources, such as from local files, dataset hubs from Hugging Face and Modelscope, and special types like arXiv, wiki, CommonCrawl, etc. data = DJDataset.load(src_file)
data = data.process(op1) # run a single operator data = data.process(op1).process(op2) # run multiple operators continuously data = data.process([op1, op2]) # run a list of operators
data.export(tgt_file) # tgt_file can be either local path or remote path Some basic interfaces of DJ-Dataset.
Moreover, following the principle of Template Method Pattern, the DJ-Dataset class establishes templated workflows for OP executing on datasets and provides several basic and unified interfaces, catering to users with diverse programming preferences, as demonstrated in Listing 5.1.1. The complexities of underlying computational engines and cumbersome runtime issues related to instantiated dataset objects are abstracted away from these interfaces, ensuring transparency and ease of use for end-users and developers alike.
New functionalities can be added by merely implementing new OP classes without altering the internal logic of DJ-Dataset. The standalone and distributed execution modes automatically recognize and switch configurations based on runtime settings. Coupled with unified function signatures, the data processing can be consolidated into a batched processing mode, not only standardizing the data interface for OP processing but also facilitating the implementation of a robust sample-level fault tolerance mechanism. These internal adaptation features will be elaborated in Sec. 5.4.2.
{listing}
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ] json ”text”: ”¡__dj__image¿Antarctica is Earth’s southernmost & least-populated ” ”continent. ¡—__dj__eoc—¿¡__dj__video¿¡__dj__audio¿ Situated almost ” ”entirely south of the Antarctic Circle, it contains the geographic ” ”South Pole. ¡—__dj__eoc—¿Antarctica is the 5th-largest continent, ” ”about 40”Most of Antarctica is covered by the ice sheet. ¡—__dj__eoc—¿”, ”images”: [ ”path/to/the/image/of/antarctica_snowfield”, ”path/to/the/image/of/europe_map” ], ”audios”: [”path/to/the/audio/of/sound_of_waves_in_Antarctic_Ocean”], ”videos”: [”path/to/the/video/of/remote_sensing_view_of_antarctica”], ”meta”: ”version”: ”0.1”, ”text_category_tags”: ”xxx” , ”stats”: ”lang”: ”en”, ”image_heights”: [336, 728], ”audio_duration”: [5], ”video_motion_score”: [1.3], Illustration of data schema of Data-Juicer 2.0.
5.1.2. Flexible Data Schema
In the context of foundation models, datasets are complex, originating from various sources and modalities. Target tasks are diverse, requiring different data formats and organizational structures. Integrating diverse data processing into a unified pipeline is challenging. Data-Juicer 2.0 adopts an extensible data representation schema that intrinsically supports multimodal data types and flexible data field customization.
An example of data sample is illustrated in Listing 5.1.1. In the data schema, each DJ-Dataset object encapsulates column-stored samples represented by a non-recursive dictionary whose level-1 fields include three categories: (1) the data payload within customizable fields such as text, images, and audios; (2) the meta information concerning this sample that stems from either its raw data or Data-Juicer’s Mappers for tagging; (3) the stats information calculated by Data-Juicer’s Filters (all Filter OPs consist of a compute_stats function, utilized for subsequent analysis and conditional filtering).
For the data payload fields, the content is primarily centered on the textual modality within the text field, capable of optionally including several text chunks split by customizable end-of-chunk tokens, which are <|__dj__eoc|> in default. Each chunk serves as a semantic unit, where the associated multimodal data within the same chunk relates to the same topic, thereby aligning with one another. Multimodal data (excluding text) are denoted by ordered customized special tokens in the text (e.g., <__dj__video>, <__dj__image>), storing entities at designated files, accessible through indexed file paths within the modality fields of samples to facilitate object sharing and lazy evaluate.
For instance, in the example shown in Listing 5.1.1, there are four chunks split by four <__dj__eoc> tokens in the text field, and two images represented by <__dj__image> tokens and stored in the images field. The former image is located in the 1st chunk and the latter image is in the 3rd chunk. All of them are visual supplementary to the textual contents in the corresponding chunks.
This token-centric representation elegantly aligns multimodal data while preserving positional information, friendly to current prevalent learning paradigms for foundation models employing next-token prediction tasks (Bai et al., 2024b; Wang et al., 2024), and beneficial for data storage—locally, data samples are represented as-is, globally, identical portions across distinct data samples are redundantly stored once.
Also note the Data-Juicer data schema’s compatibility with simple cross-modal pairing datasets (e.g., image-text pairs (Chen et al., 2015) and video-text pairs (Xu et al., 2023)), as well as complex interleaved multimodal datasets (e.g., MMC4 (Zhu et al., 2023)), owing to its chunkable and special token design. Additionally, we undertook substantial efforts to ensure Data-Juicer’s compatibility with the foundation model community, such as aligning post-tuning related fields (e.g., query, response) with the ModelScope-SWIFT repository (Zhao et al., 2024) for smooth training of 400+ foundation models. A suite of dataset conversion tools was also developed to efficiently transform various popular multimodal datasets to and from the Data-Juicer schema, enabling users to process extensive existing models and datasets within Data-Juicer. These tools also serve as demonstrations for extending support to other uncovered datasets.
5.2. Data-Juicer-Operators
5.2.1. Operator Types and Design Principles.
Operators (OPs) are fundamental units responsible for executing processing functionalities, such as enhancing the accuracy of descriptive text or filtering out images containing Not-Safe-for-Work (NSFW) content. As illustrated by the middle orange box in Fig. 1, Data-Juicer defines five previous atomic OP classes: Formatter, Filter, Mapper, Deduplicator, and Selector; alongside four new types of compositional OPs: Grouper, Aggregator, FusedOP, and ScriptOP. The first five OP classes handle dataset format conversion, sample filtering, modification, deduplication, and selection, respectively. Following the Strategy Pattern, they are designed to encapsulate diverse algorithms that can be dynamically used to process the data. Each OP has a clearly defined role and can be interchanged or modified without impacting other system parts.
{listing}
[!ht] {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]python # Default approach processes the whole dataset sequentially data.process(op1).process(op2) # or data.process([op1, op2])
# Explict fine-grained processing at the batch level data.process(FusedOP([op1, op2], bs=1000)) # Internally for data_batch in data.next(batch_size): data_batch.process(op1).process(op2) Illustration of FusedOP in Data-Juicer 2.0.
Moreover, following the Decorator Pattern, compositional OPs are provided to enhance existing functionality without modifying the prevailing OP structure while dynamically adding data processing behavior to objects. The FusedOP enables explicitly grouping multiple atomic OPs to process data in a fine-grained manner within the same data batch, opposed to the default sequential processing across datasets as demonstrated by Listing 5.2.1. The Grouper takes a DJ-Dataset as input and groups data samples into batches, which can then be input into the Aggregator for subsequent aggregation. For example, we can employ extract_entity_attribute_mapper and entity_attribute_aggregator following a key_value_grouper for meta-information extraction from textual input. Meanwhile, the ScriptOP allows users to incorporate existing Python files or execute code snippets, utilizing customized functionalities encoded within scripts like helper_func.py. This includes leveraging existing Data-Juicer command tools or integrating short Python scripts (e.g., lambda functions).
Together, these nine OP types provide robust expressive capabilities for end-to-end data processing solutions that can be embedded within a single YAML configuration file (as depicted in Fig. 4).
5.2.2. Unified Coordination of Logical Operations.
In Data-Juicer 1.0, the logical operations of different OPs were coordinated within various executors (either standalone or distributed) rather than being bound to the OPs themselves. Disentangled from the executor’s scheduling interface, it becomes challenging to determine the execution logic of different OPs, making the development and extension of individual OPs less intuitive and lacking self-explanatory qualities. In Data-Juicer 2.0, several design principles are utilized to address this issue, including the Abstract Factory Pattern, Template Method Pattern, and Single Responsibility Principle.
Specifically, a top-level OP factory class is abstracted above the aforementioned fundamental OP classes. In this class, functionalities common to all OPs are extracted, such as preprocessing of instantiated parameters, support for serialization, and configuration of OP-aware runtime parallelism. Besides, a unified run() method is implemented, maintaining a consistent interface for integration and API calls. Furthermore, parallelism in multi-processing and multi-GPU is automatically configured and decoupled from specific OPs, ensuring transparency for end users and developers, as introduced subsequently in Sec. 5.5. Lower-level OP classes define their own templated execution logic behind the run invocation. Taking the Filter class as an example, its core logic first engages the compute_stats() method to obtain statistical information based on specific metrics and then invokes its process() method to determine sample filtering based on thresholds.
On the one hand, this simplifies users’ understanding of OP types. Users can instantiate any OP and invoke it with a unified parameter signature using op.run(), thereby reducing the learning curve. The templated execution flow of various OP types is self-contained within base classes, eliminating dependencies on external executors to oversee invocation logic. On the other hand, by templating execution logic within base classes, developers can readily modify, extend, or implement new OPs. For instance, a developer aiming to customize an existing Filter does not need to rewrite a new class entirely but can inherit from a related existing class and override specific methods such as compute_stats() or others as required.
Regarding the naming and implementation of specific leaf OP classes, we adhere to extracting functionalities that are not tightly coupled with the OPs into common utility classes wherever feasible. This approach enables each OP to focus on its specific modalities and functionalities, facilitating a reduction in code complexity and enhancing clarity in understanding individual OP classes. Compared to previous implementations, the revised OP classes demonstrate easier inspection, integration, and testing. Users can utilize these robust OPs and seamlessly integrate them into their own tools or systems flexibly, both in source code or exposed RESTful API.
5.3. Processing Job Control
With the fundamental DJ-Dataset and DJ-Operators primitives established, we offer a series of control panel modules (the red box in Fig. 1) to organically combine them and accomplish data processing tasks. Executor encapsulates a series of standardized execution processes tailored for different standalone and distributed engines, leveraging modules such as Config and Monitor to accomplish end-to-end system configuration, data loading, analysis, processing iteration, data checkpointing, and more. Template workflows manage the complete data development lifecycle, including feedback, data storage, logging, and performance and operational monitoring.
The processing capability of Data-Juicer 2.0 has been encapsulated into numerous effect-proven and illustrative data recipes (in YAML format as shown in Fig. 4), catering to various vertical domains, such as multimodal contrastive data synthesis (Jiao et al., 2024) and persona-oriented data processing, maintained within the DJ cookbook online. These are user-friendly, readily deployable, and facilitate interface exposure and reuse across different levels (Sec. 4), simplifying recipe routing and tailored processing solution generation by foundation-model-based agents.
To avoid the substantial costs from trial and error for data and model development in foundation model scenarios, we further develop a Sandbox suite in Data-Juicer 2.0 for data-model co-development, serving as a specialized intermediate layer connecting data processing jobs to numerous open-source infrastructures of model training and evaluation. The suite offers template workflows that extend beyond dataset-only development by incorporating cost-effective model training and evaluation signals, and quantitatively studying interplays among data, model, and compute. Users can easily conduct small, quick, and comparative experiments to find insights and superior data recipes, which can then be scaled to larger models and datasets, thereby optimizing the return on investment in data-model co-development. Additional details and empirically validated support can be found in (Chen et al., 2024c).
5.4. Internal Adaptation for DJ-Dataset
5.4.1. Engine-Agnostic Processing
In Data-Juicer 1.0, the data processing pipeline is implemented using two distinctive execution modes: a standalone mode tailored for single-node operations and a distributed mode designed for scalability over multiple nodes. These modes exploit different dataset classes, each with a unique set of functionalities and interfaces. The default standalone execution mode employs the Hugging Face Dataset class (Lhoest et al., 2021), which provides a rich array of encapsulated functions, such as dataset.map() and dataset.filter(). It’s also equipped with configurable batch processing capabilities essential for various computational needs. In contrast, the distributed execution mode leverages the Ray Dataset class (Moritz et al., 2018), which scales effectively across multiple nodes.
Despite both dataset classes using a storage format based on Apache Arrow (Arrow, 2023), they exhibit significant differences in the behaviors and interfaces exposed. For instance, the Ray Dataset delineates individual and batched sample processing using separate methods: map and map_batches. It also allows to specify GPU counts for optimized scheduling. Meanwhile, the Hugging Face Dataset excels in supporting a broad spectrum of data modalities, such as image and audio. Moreover, the Hugging Face Dataset is typically applied in read-heavy data processing scenarios, such as in-memory tokenization and tensor reshaping, which are crucial for training deep learning models. For processing tasks in the context of foundation models, especially those involving synthesis operations, write-heavy procedures warrant attention. Here, the Ray Dataset provides flexible data exportation techniques advantageous to such tasks.
As highlighted in Sec. 5.1, we introduce a top-level DJ-Dataset class in Data-Juicer 2.0, along with common functions to bridge the variety of interfaces and implementation specifics across diverse computational engines. Besides the support of Hugging Face Dataset and Ray Dataset, our DJ-Dataset is also seamlessly extendable to support distributed computing within Alibaba Cloud ecosystems, thanks to the compatibility between ours and MaxFrame-DataFrame classes with intermediate in-memory formats like Pandas (pandas development team, 2020) and NumPy (Harris et al., 2020) or external-memory formats like Parquet. Thus, we develop unified wrappers that extract the core execution functions of different DJ-Operators into User-Defined Functions for the MaxCompute engine. Compared to Spark, MaxCompute is compatible in both syntax and runtime on Alibaba Cloud nodes and can be considered a commercially optimized version. An internal empirical comparison shows that MaxCompute SQL achieves 50% better performance than native Spark SQL.
The Deduplicators within Data-Juicer 2.0. Fuzzy deduplication is complex, involving a mixture of operations such as map, filter, group by, aggregate, and join (Broder et al., 2000). The support and performance of these operations vary across different engines, especially in large-scale scenarios. To demonstrate the engine-agnostic feature of Data-Juicer 2.0, we use our minhash_deduplicator as an example, which supports the aforementioned three different engines. Users only need to specify algorithm-specific parameters such as jaccard_threshold and num_permutations. These parameters remain consistent regardless of the engine used, while engine-specific details and optimizations are completely transparent to the user. For instance, we utilize Ray Actors to implement our Ray-based deduplicator, starting with a load-balanced, distributed union-find algorithm (Kim et al., 2024) and introducing a hash-based aggregation process to enhance memory utilization and efficiency. This method avoids fragmented unions caused by Ray’s native groupby operation, which is computationally expensive in typical LSH implementations with traditional big-data engines (Mou, 2023). As a result, we achieve a 3.3x speedup over our vanilla Ray version. More performance evaluations are presented in Sec. 6.1.4.
5.4.2. Fault Tolerance
In practical processing scenarios, datasets often contain schema-incompatible or corrupted data elements, such as improperly formatted JSON objects or damaged images that cannot be read. This issue becomes increasingly important with large-scale datasets, as processing tasks may extend over several hours or even days. In Data-Juicer 1.0, corruption of a single sample would halt the entire processing task, resulting in a waste of computational resources and the loss of already processed samples.
To address this issue, Data-Juicer 2.0 introduces a sample-level fault tolerance mechanism designed to enhance processing reliability by providing a worst-case guarantee. A unified exception manager is implemented to automatically capture runtime errors with customizable handlers during the processing of each OP. By default, dataset processing operations are performed at the data-batch level for a general handler. As a result, the system can easily bypass problematic samples by skipping the affected batch (as demonstrated in Fig. 5), while actively tracking and reporting these cases for subsequent debugging and correction. This ensures a seamless user experience and minimizes retry costs in scenarios involving large-scale data processing. Users have the flexibility to either discard these samples or mark them for future reprocessing.
Figure 5. Illustration of fault tolerance for “bad” data.
Of particular note is the fact that implementing this mechanism in Data-Juicer 1.0 presented challenges due to the rigid schema consistency mandates imposed by the underlying Hugging Face and Ray dataset classes. The newly refactored batching interfaces facilitate the unified construction of compatible empty samples based on schemas accessed during exception handling, thereby enabling seamless integration with valid dataset entries.
5.4.3. Streaming I/O and Subset Splitting
Memory constraints frequently emerge as a bottleneck in data processing tasks associated with foundation models. Memory demands can be substantial, and its precise usage is difficult to predict beforehand. Influential factors include the actual storage demands of objects such as text, image, and audio within individual dataset samples, as well as the storage demands of auxiliary or newly generated objects at runtime. These can stem from varying OP specifications, model sizes, synthesis data volumes, intermediate variables, the number of concurrent processes, and specific computational engines.
To effectively adapt to a diverse range of data processing scenarios with varying data volumes and available computational resources, we introduce streaming loading and data pre-splitting capabilities in Data-Juicer 2.0. They collectively facilitate improved computational resource utilization, and provide potential for a flexible programming space of hybrid stream and batch processing.
Firstly, Data-Juicer 2.0 offers a streaming loading interface, addressing the current lack of native support in the Arrow framework underlying Hugging Face and Ray Datasets for streaming JSONL data. As many foundation model datasets use JSONL, we developed an in-house patch to alleviate Out-of-Memory (OOM) issues.
Secondly, we develop a user-friendly script that automatically pre-splits the original dataset based on two observations: (1) The size limit of the underlying Apache-Arrow block, and (2) The inherent automatic block-splitting strategy in Ray. When there are tens of thousands of nodes but with only a few dataset files, Ray would split the dataset files according to the available resources and distribute the blocks of the dataset to all nodes, which brings a huge network communication cost and decreases the CPU utilization of each node. Thus, we split the original dataset into smaller 128MB files in advance automatically according to the dataset size and the number of distributed nodes, trying to adapt the features of Arrow and Ray for better performance (an empirical study can be found in Sec. 6.1.4). This approach reduces location and reprocessing costs associated with fault tolerance and helps mitigate network exchange overheads, especially beneficial in contexts involving large-scale multimodal data, as well as in scenarios that require handling global objects of Ray Actor in distributed modes.
5.5. Internal Adaptation for DJ-Operators
This section explores the internal adaptations developed for DJ-Operators, which are crucial for optimizing resource allocation and user experience without requiring users to understand hardware specifics or implementation details. We implement several automatic adaptation features for resource management, aiming at balancing resource constraints and operational efficiency within Data-Juicer 2.0. These strategies include automatic OP reordering at the recipe level (Sec. 5.5.1), automatic parallelism at the OP level (Sec. 5.5.2), and automatic data insight mining to assess the impact of each OP on data samples, considering both upstream and downstream OPs (Sec. 5.5.3). These features are encapsulated in a dedicated Adapter class, which uses a probe_small_batch() method to systematically probe and analyze essential information by applying individual OPs on randomly sampled data in runtime, with default sample size as min(1000, remaining_data_size).
5.5.1. Workloads-Aware OP Reordering
At the recipe level, we introduce a new probe-based OP reordering strategy. In Data-Juicer 1.0, an OP fusion optimization was proposed to eliminate redundant computations for the Filter OPs, which involved three key steps: detection and grouping of fusible OPs, OP fusion, and OP reordering. The reordering strategy aimed to position more time-consuming OPs at the end of the group to process fewer samples to save time, as some were filtered out by preceding OPs. It is assumed that the fused OPs are the most time-consuming, and only the fused OPs are moved to the end of each group.
However, the assumption is not always correct and the prior reordering strategy omit the unfused OPs, rendering it greedy and suboptimal, especially when applied to diverse datasets and data recipes characterized by varied data distribution and OP orchestration. In Data-Juicer 2.0, we advance the reordering strategy to an adaptive approach, which is workload-aware and can be applied automatically to unfused OPs as well.
Specifically, before processing the full dataset, Data-Juicer 2.0 utilizes the probe functionality of Adapter to obtain estimated processing speeds for individual OPs relevant to specific input datasets. When processing the entire dataset begins, OPs in each group (including the unfused ones) are reordered based on the probed speeds and the commutativity of the Filter OPs. For the fused OP, assuming that there are $n$ fusible OPs in it and their probed speeds are $v_{i}$ where $i\in\{1,2,\cdots,n\}$, the estimated speed of the fused OP can be calculated as:
(1) $v_{fused}=\frac{N}{T_{total}}=\frac{N}{\sum_{i=1}^{n}\frac{N}{v_{i}}}=\frac{1}{\sum_{i=1}^{n}\frac{1}{v_{i}}},$
where $N$ is the total number of samples to be processed by the fused OP and $T_{total}$ is the total time cost of it. Then, faster OPs are prioritized, while slower ones are deferred to later stages. This probe-based approach identifies the globally optimal reordering solution for each OP group, outperforming the suboptimal strategies of the previous version, as empirically validated in Sec. 6.2.1.
5.5.2. Automatic Operator-Wise Parallelism
In data processing for foundation models, it is crucial to recognize that different OPs require vastly different computational resources. Model-based OPs often need several gigabytes of GPU memory, while simple rule-based OPs may only need minimal CPU processing. Therefore, using a uniform parallelism granularity across all OPs in a data pipeline can cause OOM issues for some and resource underutilization for others. To address this challenge, we introduce several automatic mechanisms for OP-wise parallelism.
  • Model-based OPs that integrate substantial models require considerable computational time, potentially spanning hundreds of hours on CPUs for large datasets. To mitigate this, Data-Juicer 2.0 expedites these OPs by automatically leveraging CUDA and GPUs when available. Given the diverse memory requirements of large models, we utilize the Adapter component to conduct quick VRAM benchmarking prior to full-size dataset processing. This information is systematically assigned to the mem_required parameter of the respective OPs. To further alleviate GPU memory demand, we also integrate quantization libraries such as vLLM (Kwon et al., 2023) to enable efficient inference of leveraged models.
  • For non-model-based OPs, attributes such as cpu_required and mem_required are crucial. In Data-Juicer 2.0, users can specify a global parallelism level, typically aligning with the available processor count. Meanwhile, we calculate the adaptive number of processors to finely optimize the entire processing pipeline, aiming to maximize resource utilization (90% by default) as much as possible. For this purpose, the Adapter is also employed dynamically and instrumental in determining precise cpu_required and mem_required values at runtime.
  • For general-purpose and I/O-intensive OPs, Data-Juicer 2.0 enables batched processing, using a robust default batch_size parameter guided by performance saturation trends shown in Sec. 6.2.3. Batched processing reduces I/O overhead, boosts efficiency per processor, and enhances overall speed. Additionally, we introduce hierarchical parallelism for OPs involving multimodal data I/O, such as image_aspect_ratio_filter. These OPs utilize multi-process and GPU parallelism, along with multiple threads, to handle data batches more efficiently, taking the concurrent opportunities between I/O and computation latencies.
5.5.3. Automatic Operator-Wise Insight Mining
The cumulative impact of a sequence of OPs on a dataset can be complex, and the combined effectiveness is not always greater than that of a single OP. This is demonstrated by experiments on cutting-edge foundation models for text and video (Chen et al., 2024c). Traditionally, systems like Data-Juicer 1.0 (Chen et al., 2024a) and Falcon (Penedo et al., 2023) primarily generated coarse-grained visualizations such as Sankey diagrams to show percentage changes in data volume, mainly through Filters.
Figure 6. Illustration of OP-wise insight mining.
To formulate better data recipes, we introduce fine-grained OP-level insight mining and visualization, extending beyond Sankey diagrams. As shown in Fig. 6, our system provides comprehensive tracking of data statistics (e.g., stats.perplexity) and category tags (e.g., meta.text_category_tags). After the processing completion of each OP, we use the Adapter’s probing function along with Data-Juicer’s embedded Analyzer. Built-in category tagging OPs across modalities and analytic Filters help obtain histograms of interested semantic categories and relevant statistics.
Data-Juicer 2.0 uses these analytics to automatically assess changes in metrics between successive OPs, generating detailed visualization reports that highlight significant differences along lineages. This reporting aids in analyzing each OP’s significance, understanding correlations, and optimizing data recipes. For instance, if an image-text matching Filter (using a model like BLIP-2 (Li et al., 2023a)) reduces text length statistics significantly, it may indicate noisy, too-long captions that could affect textual description accuracy for text-image pairs.
6. System Experiments
In this section, we present practical patterns and anti-patterns of Data-Juicer 2.0 based on extensive performance evaluations.
6.1. Cloud-scale Data Processing Efficiency
6.1.1. Experiment Settings.
We start with a base dataset comprising 560k image-text pairs (about totally 5.6 million textual tokens). We expand this dataset by factors of {2, 4, 20, 100, 500, 2500, 12500, 125000}, resulting in nine datasets scaling to 70B data samples. These datasets are categorized into three scales: small (1x, 2x, 4x), medium (4x, 20x, 100x), and large (100x, 500x, 2500x, 12500x, 125000x). For different scales, we prepare both multimodal and text-only data processing recipes, each containing five OPs. We run the recipes on these datasets using various computing engines (Standalone, Ray, MaxCompute) with different Alibaba Cloud resources, including ECS instances, PAI-DLC, and MaxCompute, spanning 1 to 100 nodes and 64 to 12,800 CPU cores. To ensure a fair comparison, we use the same worker node configuration, such as CPU frequency, across different computing engines.
(a) Small-scale multimodal datasets.
(b) Small-scale text-only datasets.
(c) Medium-scale multimodal datasets.
(d) Medium-scale text-only datasets.
(e) Large-scale text-only datasets.
(f) Resource utilization comparison.
Figure 7. Data processing time comparison with different computing engines, dataset sizes, and workloads.
6.1.2. The Case of Small Scales
As shown in Fig. 7a, when processing small-scale multimodal datasets, the standalone mode with Hugging Face Dataset is efficient and comparable to the Ray mode with a single node. Additional Ray nodes provide further but limited acceleration (speedup ratios between 138% and 226% with 4 nodes). For text-only datasets (Fig. 7b), the standalone mode remains efficient. However, for the Ray mode, 4x node increments yield smaller speedups (148%) or even increased processing times due to dominated I/O and communication costs. Thus, with Data-Juicer, processing datasets with hundreds of thousands of samples on a single machine is both efficient and cost-effective for most users.
6.1.3. The Case of Medium Scales.
When datasets are scaled to 56M samples, processing times increase significantly with the standalone mode, thus, its performance is omitted in Figures 7c and 7d. Here, the Ray mode outperforms in all instances, demonstrating considerable speedups with increasing node counts, making it the recommended choice for medium-scale scenarios. Moreover, compared to native ECS nodes (green lines), the Ray mode on PAI-DLC (red lines) is faster, saving 24.8% of processing time due to the Alibaba Cloud’s optimization dedicated to cluster networking.
6.1.4. The Case of Large Scales.
For datasets at the 70B-sample scale, professional cloud-scale distributed data processing products are advantageous. Users benefit from the vast cloud computing resources without dealing with the intricacies of setup and management. From our experiments, we recommend Ray-DLC for multimodal recipes and MaxCompute for text-only recipes at this scale. (1) For multimodal recipes, using 3200 Ray-DLC cores process datasets in 1780.86s and 7083.5s for 500x and 2500x dataset sizes, respectively, indicating good scalability. On the other hand, the MaxCompute engine requires 1.5 times more processing time using the same resources, due to the challenges of loading large-size multimodal data. (2) For text-only recipes (Fig. 7e), although the Ray mode benefits from additional cores even at the scale of ten thousand, MaxCompute is the fastest, requiring about 1/4 of the time and using 1/2 fewer cores. This is attributed to MaxCompute’s co-optimization of distributed computing and storage, a feature not as advanced in Ray’s implementation.
Impact of I/O and Data Splitting. I/O plays a crucial role in data processing workloads. For distributed processing, large-scale datasets need to be distributed to nodes, stored on shared disks (e.g., NAS), or on cloud storage services (e.g., Alibaba Cloud OSS and CPFS). All these introduce unavoidable I/O and network communication costs. In our experiments shown in Fig. 7, we use Alibaba Cloud standard CPFS, finding that NAS and OSS can increase time costs by 20-30% compared to CPFS. Furthermore, with another AI-oriented CPFS product that has a 3x larger network bandwidth and optimized RDMA, processing the 12500x dataset with 3200 cores in Ray mode takes 1625s, 2.7 times faster than the standard CPFS version (4396s). We then scale up the dataset to 125000x, resulting times as 14617s and 7611s with 3200 and 6400 cores respectively, further demonstrating the system’s efficiency and scalability.
Additionally, the adaptive subset splitting capability for Ray mode, introduced in Sec. 5.4.3, offers 2$\sim$3x speedups in our experiments. For example, with 100 Ray-DLC nodes and 12800 cores, our subset splitting method reduces the processing time of 12500x dataset from over 5000s to about 2000s. Analysis of Ray’s resource utilization shows that its automatic block-splitting strategy incurs high network communication costs and low CPU utilization when many nodes handle a few dataset files. Our strategy pre-splits datasets according to size and node count, optimizing alignment with Ray’s features and enhancing efficiency. As seen in Fig. 7f, this strategy reduces peak network I/O from about 160MB/s to 60MB/s and achieves more consistent CPU utilization across all nodes, thus accelerating dataset processing.
Table 1. Deduplication time (minutes) with Ray nodes across varying dataset sizes and CPU counts (160 cores per node).
{tblr}
width = colspec = Q[154]Q[240]Q[240]Q[240], hline1,4 = -0.08em, hline2 = -, # CPU & 200GB Time 1TB Time 5TB Time
4*160 11.13 min 50.83 min 285.43 min
8*160 7.47 min 30.08 min 168.10 min
Large-scale Deduplication. We tested the MinHash-based RayDeduplicator with datasets from Redpajama (Computer, 2023) and Common Crawl chunks (Crawl, 2023), sized at 200GB, 1TB, and 5TB, using CPU counts ranging from 640 to 1280 cores. As shown in Table 1, RayDeduplicator efficiently scales with increased data size and computing resources, indicating its capability to handle large-scale deduplication tasks effectively. When the data size increases by 5x, the processing time increases by 4.02$\sim$5.62x. When the number of CPU cores doubles, the processing time decreases to 58.9$\sim$67.1of the original time. Notably, RayDeduplicator can process 5TB of data in 2.8 hours using 8 Ray nodes (8*160 CPU cores). In contrast, NVIDIA’s NeMo Curator, which leverages cuDF (cuDF Team, 2024) and Dask (Team, 2024a), takes 1.8 hours to process 1.1TB of Redpajama data using 64 A100 GPUs (64*6192 CUDA cores), as shown in (Team, 2024b).
6.2. Ablation Study on Runtime Adaptations
6.2.1. Automatic Workloads-aware OP Reordering.
We evaluate the performance improvements due to our probe-based OP reordering on two recipes with different numbers of OPs. The simple recipe contains 5 OPs, with 2 fusible OPs, while the complex recipe includes 13 OPs, with 5 fusible OPs. We run these recipes on the base dataset used before (580k image-text pairs), comparing processing times for all OPs, fusible OPs, and other OPs.
Fig. 8 illustrates the results. Generally, OP fusion is effective for accelerating data processing, and automatic OP reordering offers further improvements based on OP fusion. Notably, OP reordering is more effective in complex data recipes (the right sub-figure), saving more time, especially for fusible OPs (46.09% v.s. 70.22%).
Figure 8. Processing time for simple (left) and complex (right) recipes. Processing times for different types of OPs are scaled according to the time of baseline without OP fusion.
6.2.2. Automatic GPU Resource Allocation
Efficient model inference on GPUs is critical for the processing speeds of model-based OPs. We select four image-related OPs, including an image-only OP and three image-text cross-modality OPs, integrating various models that require different GPU memories. Using a test dataset of 10k image-text pairs, we process these OPs on a machine with 64 processors and compare the times with and without GPU support in Data-Juicer 2.0. Table 2 shows these comparative results.
Table 2. Multimodal OPs processing time comparison between CPU cores and GPUs on 10k samples. “np” denotes the number of processors.
{tblr}
width = colspec = Q[520]Q[100]Q[40]Q[150]Q[150], cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=2, hline1,7 = -0.08em, hline3 = -, hline2 = 4-5, Multimodal OPs & VRAM np Processing Time
CPU GPU
image_captioning_mapper ~16GB 32 8668.87s 305.44 s
image_diffusion_mapper ~8GB 64 100 h ~1 h
image_text_similarity_filter ~1.5GB 64 73.04 s 35.84 s
image_nsfw_filter ~1GB 64 102.59 s 39.74 s
As the results show, using GPUs saves at least 50% of processing time for all selected model-based OPs, especially for large, slow models like BLIP-2 (Li et al., 2023a) used in image_captioning_mapper and SDXL model (Rombach et al., 2022) in image_diffusion_mapper. Compared to the CPU version, GPU usage and adaptive resource allocation are much more efficient and necessary. Due to the GPU memory limit of around 80GB, we can only allocate at most 4 models of image_captioning_mapper OP on a single GPU, so the number of processors for this OP is automatically reduced to 32, preventing OOM errors.
6.2.3. Batched Data Processing
We compare the efficiency of single-sample processing in the previous Data-Juicer 1.0 and batched processing in Data-Juicer 2.0. We use the previous 560k multimodal dataset, processed by 16 processors. Each processor handles about 30k samples. We select four recipes representing different scenarios: Filter-Heavy, Mapper-Heavy, Text-Heavy, and Multimodal-Heavy. Each recipe contains four OPs of the specified “heavy” type and one of another type. We run these recipes with five different batch sizes and summarize the results in Fig. 9a. From these experiments, we conclude: (1) Batched processing is always more efficient. Larger batch sizes consistently speed up data processing in all scenarios, reducing processing time by 84%. (2) Efficiency gains from larger batch sizes plateau beyond a certain threshold. Specifically, batch sizes of 100 or more yield similar benefits. (3) A batch size of 1000 is recommended, generally showing the most efficient processing in our experiments. Consequently, 1000 is selected as the default batch size in Data-Juicer 2.0.
(a) Different batch sizes and recipes.
(b) Different parallelism levels.
Figure 9. Data processing efficiency of OP-wise hierarchical parallelism. “bs” denotes the batch size, and “nt” dennotes the number of threads.
6.2.4. Automatic OP-wise Hierarchical Parallelism
To verify the efficiency of the automatic OP-wise hierarchical parallel strategy, we select the multimodal OP image_aspect_ratio_filter as example. We conduct experiments with various processor counts, batch sizes, and thread numbers, covering three parallel levels: multiprocessing, batched processing, and multithreading. This OP processes approximately 560k image-text pairs. Fig. 9b demonstrates the time consumption and efficiency gains.
As depicted, significant speed improvements are consistently achieved across all parallel levels. Beyond the multiprocessing and batched processing strategies examined earlier, multithreading further reduces processing time. This OP benefits from multithreading due to its intensive I/O procedures (reading images), which balance I/O and CPU utilization. Most other multimodal OPs also require heavy data read/write operations, thus benefiting from the multithreading strategy.
7. Conclusions
Data-Juicer 2.0 emerges as a versatile scaffold in the evolution of foundation models, providing efficient and adaptive data processing solutions for handling the diversity and scale of modern datasets. Our re-envisioned architecture leverages multi-layered adaptability to coordinate different modules, operators, and runtime environments. Extensive evaluations reveal its high performance across diverse cloud-scale workloads. By open-sourcing this system, we aim to foster a vibrant community of contributors and users, encouraging collaborative development and driving innovation to underpin the next generation of foundation models. Future works include further enhancing the system in terms of adaptability using model-driven agents (Gao et al., 2024), scalability by optimizing the single-node transmission bottleneck in Ray’s head node (Team, 2024d), and efficiency by supporting GPU backend engines like NeMo Curator (Jennings et al., [n.d.]).