### v2

# Data-Juicer 2.0: Cloud-Scale Adaptive Data Processing for and with Foundation Models

Abstract
The burgeoning field of foundation models necessitates advanced data processing mechanisms capable of harnessing vast and valuable data with various types used by these models. Nevertheless, the current landscape presents unique challenges that traditional data processing frameworks struggle to handle effectively, particularly in handling the complexity of multimodal data. In response, we present Data-Juicer 2.0, a data processing system backed by 100+ data processing operators spanning text, image, video, and audio modalities, supporting more critical tasks including data analysis, synthesis, annotation, and foundation model post-training. With seamless compatibility and dedicated optimization for popular dataset hubs like Hugging Face and computing engines like Ray, it improves upon its predecessor in terms of usability, efficiency, and programmability. It features an easily accessible user interface layer that supports decoupled Python interactions, RESTful APIs, and conversational commands. It contains a new runtime layer optimized for adaptive execution and management across varying dataset scales, processing demands, and computational environments, while hiding unnecessary system details. Extensive empirical evaluations demonstrate Data-Juicer 2.0’s remarkable performance and scalability, highlighting its capability to efficiently process TB-level data with 10k+ CPU cores. The system is publicly available and has been widely adopted in diverse research fields and real-world products such as Alibaba Cloud PAI. We actively maintain it and share insights from practical feedback, with the goal of facilitating research and application of next-generation foundation models.
11footnotetext: Co-first authors.22footnotetext: Equal contribution.33footnotetext: Corresponding authors, email addresses: {yaliang.li, bolin.ding}@alibaba-inc.com
1 Introduction
Data Processing Challenges for Foundation Models. Modern foundation models require sophisticated data pipelines to handle multimodal inputs and evolving paradigms for pre-training and post-training. While existing frameworks address specific aspects of text processing together2023redpajama ; DolmaToolkit or traditional big data workloads apache-spark , they lack essential capabilities for contemporary multimodal workflows. Three critical gaps emerge:
Multimodal Processing Limitations: Current tools provide inadequate support for cross-modal alignment and semantic-aware transformations crucial for vision-language models li2024mgm . The transition from text-only systems like the inaugural version, Data-Juicer 1.0.  [^1] In this paper, references to Data-Juicer 1.0 pertain to the SIGMOD publication chen2024dj and the open-source codes versioned with the v0.1.2 Based on it, there are over 100 pull requests (PRs) for updating to Data-Juicer 2.0, including about 40k and 8k lines of Python code addition and deletion, respectively. to multimodal processing introduces architectural challenges in handling heterogeneous data types and inter-modal relationships.
Efficiency-Scalability Tradeoffs: Traditional big data engines dask and Python-based solutions lhoest-etal-2021-datasets struggle with foundation models’ unique computational patterns - simple per-sample operations applied at petabyte scales. This creates reliability risks in large-scale processing scenarios where late-stage errors can invalidate terabytes of computation.
Ecosystem Fragmentation: Disjoint APIs across popular frameworks (Hugging Face Datasets lhoest-etal-2021-datasets , Ray Data ray ) force practitioners into suboptimal workflow choices. The lack of unified abstractions hampers portable optimization and cross-platform execution.
Architecture Overview. Data-Juicer 2.0 addresses these challenges with a layered architecture optimized for foundation models (Fig. 1), enabling multimodal data processing at scale:
Capability Layer: Extends Data-Juicer 1.0’s 50 text-only operators (OPs) for pre-training to 150+ multimodal OPs supporting text/image/video/audio processing and more post-training tasks. New OPs integrate foundation models (e.g., Tongyi-Qwen qwen3 , SDXL podell2023sdxl ) for semantic-aware filtering and cross-modal synthesis while maintaining compatibility with mainstream dataset hubs hf-dataset-hub ; ms-dataset-hub .
Interface Layer: Provides multi-level APIs balancing flexibility and accessibility. Low-level Python APIs enable custom workflows, while RESTful endpoints and a visual editor support rapid prototyping. Novel agent integration allows natural language specification of data pipelines.
Runtime Layer: Introduces four key enhancements: (1) A unified Data-Juicer-Dataset abstraction spanning local to cloud-scale execution environments; (2) Decoupled OPs with automatic optimization through adaptive batching and resource allocation; (3) Control plane supporting fault-tolerant execution and data-model co-development chen2024djsandbox for insight mining; and (4) Adapter components enabling automatic hardware-software configuration and parallelism across diverse deployments.
Contributions. Our contributions are summarized below:
Enhanced System and Adaptive Techniques. Incorporating feedback from numerous users and cutting-edge applications built upon its predecessor, Data-Juicer 2.0 emerges as a new open-source system with enhanced multimodal processing capabilities. It emphasizes dedicated multi-layered adaptability and system optimization, integrating cloud-scale practices and diverse computation engines to dynamically and efficiently meet data processing demands.
Extensive Evaluation and Usages. We conduct thorough experimental analyses to assess the system’s data processing capabilities under varied workloads, such as filtering-intensive, model-based, and semantically editable multimodal scenarios, using datasets ranging from millions to tens of billions of samples. We provide actionable insights and performance trade-offs across different use cases and resources, involving Ray and Alibaba MaxCompute maxcompute2024 with up to 100 nodes and 12,800 cores.
Community and Applications. We have open-sourced this new system at https://github.com/modelscope/data-juicer, fostering sustained maintenance and engagement through practical events qin2024synergy ; dj-competition , such as tutorials, surveys, data competitions, and co-optimization with community like Apache Arrow, Ray, and NVIDIA-Nemo-Curator teams.  [^2] Related PRs and docs can be exemplified in links for Arrow, Ray and Nemo respectively. Data-Juicer 2.0 also facilitates many foundation model researches such as those from Alibaba Tongyi ge2024data ; qin2024federated ; bai2024federated ; jiao2024img ; ling2025diversity ; xu2025mindgym ; zhou2024humanvbench , and serves as the operator base for multiple Alibaba Cloud products such as PAI Designer paidesigner2024 , indirectly benefiting hundreds of internal and external customers across various real-world AI businesses. One of them has been running stably for over five months, processing data at a scale exceeding terabytes.
Figure 1: The overview of Data-Juicer 2.0.
2 Processing Capabilities Beyond Text & Pretraining
Data-Juicer 2.0 extends its predecessor, Data-Juicer 1.0—which focused on pretraining tasks for text-only LLMs—to support the processing of multimodal datasets on more tasks for a wide range of foundation models with about 100 new OPs. Images, videos, and audios are all valid inputs of Data-Juicer 2.0, and it’s able to process datasets for post-training tasks, including supervised fine-tuning (SFT), Reinforcement fine-tuning (RFT), and so on. A detailed list of these OPs is placed in Appendix B. For clarity and ease of reference, the OPs are categorized from the following aspects, and their distributions are visually displayed in Fig. 2:
(a) Operator type.
(b) Modality type.
(c) Function type.
(d) Implementation type.
Figure 2: Distribution of new OPs across various dimensions.
Operator Types: These new OPs are built on both original types in Data-Juicer 1.0 and types newly introduced by Data-Juicer 2.0. Among them, Formatters load datasets; Mappers edit samples; Filters compute data stats and remove samples accordingly; Deduplicators find redundant samples; Selectors sample data based on preferred ranks or rules; Groupers batch samples, and then Aggregators combine them into one. Notably, Data-Juicer 2.0 introduces a new type of OP named HumanOP. It’s built upon Label Studio LabelStudio and involves asynchronous human annotations and feedback during data processing and time-delayed downstream training tasks. This new OP helps to build SOTA human-in-the-loop procedures, such as reinforcement learning from human feedback (RLHF). These OP types will be detailed later in Sec. 4.2. Statistically, about 90% of OPs are Mappers and Filters, reflecting trends of user needs in the foundation model domain. New variants in other types, such as Formatters and Deduplicators, remain stable due to the broad applicability of already supported formats (e.g., JSONL, Parquet, MP4) and classical algorithms like MinHash broder1997resemblance .
Modality Types: The majority of new OPs are concentrated on video/image/audio/text-only processing, with about 20 OPs dedicated to cross-modal data processing. Among these new OPs, almost 3/4 of them are aimed at multimodal data processing, covering images, videos, and audio. There are both single-modality OPs, such as video_motion_score_filter that scores the dynamics of videos, and cross-modal OPs such as phrase_grounding_recall_filter and video_captioning_from_summarizer_mapper, which measure the alignment between different modalities and generate contents from one modality to another, respectively. Detailed showcases are in Appendix C.
Function Types: Data cleaning and analysis operations constitute nearly 1/3 of the new OP suite. In particular, Data-Juicer 2.0 introduces about 50 OPs for data synthesis and augmentation on cross-modal, post-training, and reinforcement learning scenarios, enabling any-to-any generation among 4 supported modalities and improving textual dialogs based on varied information and needs. Furthermore, new privacy protection OPs, such as removing not-safe-for-work (NSFW) contents or blurring human faces, and several wrapper OPs for established 3rd-party tools, such as FFmpeg ffmpeg , have been integrated, enabling users to conveniently invoke existing professional commands.
Figure 3: Distribution of fine-grained function types for text-only OPs.
Besides, 25 text-only OPs are introduced for post-training tasks. We tag them with fine-grained function types and summarize their distribution in Fig. 3. Among them, there are information extraction and synthesis OPs for curating dialog samples, as well as calibration and refinement OPs for optimizing questions and responses in SFT and RFT tasks. Beyond previous rule-based OPs to assess quality, difficulty, and diversity of datasets in Data-Juicer 1.0, more foundation-models-based OPs emerge to score the datasets from diverse dimensions and analyze the rationale of given scores. This practice is flexibly and useful for vertical and cross-task scenarios. For example, rule-based methods usually struggle to analyze the difficulties of math and financial problems, while this kind of OPs generally perform better with reduced customization and tuning efforts.
Implementation Types: The new OPs include both novel algorithms from Data-Juicer 2.0 and implementations based on state-of-the-art (SOTA) methodologies schuhmann2021laion ; zhu2023multimodal ; gadre2024datacomp from the community. Some OPs offer multiple versions to accommodate varying computational resources, such as CPU-only (e.g., with OpenCV opencv_library ) or GPU-based configurations (e.g., with RAFT teed2020raft ). In the distribution, model-based OPs dominate, as semantic-aware processing often requires advanced models, such as large models for general-purpose understanding and generation via SDXL, GPT, and Qwen. Data processing with foundation models is becoming more popular and helpful.
3 Towards a More Accessible System
Data-Juicer 2.0 involves substantial work to improve the system accessibility over Data-Juicer 1.0, including a variety of user interfaces and a smooth experience for both beginners and light users.
On the one hand, in terms of user interfaces, in addition to the way to process data with an all-in-one configuration file provided in the previous version, we introduce 4 more user interaction methods. (1) Low-level APIs. We make the underlying implementation more transparent, and expose many extensible interfaces of Data-Juicer-Dataset, Data-Juicer-Operators, such that users can conveniently integrate Data-Juicer in their code. (2) RESTful APIs. For some web server usages, we provide one-click generation of high-performance web APIs capable of automatically discovering, registering, and adapting OP classes and tools. Users can start the server easily and trigger data processing across computing nodes. (3) Web tools. Based on RESTful APIs, Data-Juicer’s capabilities are integrated into Alibaba Cloud’s visual modeling product, PAI-Designer paidesigner2024 , which provides a drag-and-drop UI for users to organize the data processing pipelines and allows users to make use of more products in the Alibaba Cloud ecosystem. (4) Natural language interaction. It’s challenging to know when and how to use each of the 100+ diverse OPs. In response, we adopted AgentScope gao2024agentscope , a multi-agent platform, for low-code integration, using prompt-optimized reasoning and acting (ReAct) Yao-arxiv-2022-ReAct agents to align OP functionalities with our RESTful APIs. Users only need to tell the agents how they want to process the dataset, and the system will automatically handle the job analysis and execution. More details about these interfaces are in Appendix D.
On the other hand, as the 100+ OPs in Data-Juicer 2.0 cover quite diverse multimodal data and training tasks, the complete capabilities require heavy dependencies and take the risk of runtime errors. This may slow down the system installation, initialization, and troubleshooting. We thus prepare a minimal set of requirements and split OPs’ dependencies into subgroups based on their modalities and usage categories. Users only need to install the lite Data-Juicer in a quick way, and the optional dependencies will be lazy-loaded when using specific OPs. Moreover, we maintain an automated unit and regression testing mechanism, ensuring over 85% test coverage. These actions make our system more user-friendly, helping both beginners and light users get started with it.
4 Towards a More Flexible, Robust, & Efficient Runtime
4.1 Data-Juicer-Dataset
Unified Execution Abstraction. Data-Juicer 2.0 introduces a Data-Juicer-Dataset class that abstracts heterogeneous computational engines (Hugging Face Dataset, Ray Data, MaxFrame-DataFrame), while preserving native API compatibility through a Facade-pattern design gamma1995design . It provides unified interfaces for standalone/distributed execution modes, enabling transitions between environments while hiding engine-specific complexities. The class supports chainable processing workflows via templated methods with multiple process() call manners. More implementation details about these modes are provided in Appendix E.1 and Listing E.1.
Reliable Data Loading. Data loading is refactored to a more systematic way instead of only specifying the dataset path. The new DatasetBuilder class supports explicit dataset source specification (e.g., local, remote) and customizable configuration for loading datasets. To improve data loading reliability, a novel DataValidator module is introduced to check and validate data sources, schemas, and whether the dataset meets the goal of data processing. For example, post-tuning data should contain dialogs, and image paths should exist in image-captioning datasets.
Token-Aligned Data Schema. Our intermediate schema represents multimodal data through special tokens (e.g., <__dj__image>, where “dj” stands for Data-Juicer) in text fields, with chunk-based alignment using <|__dj__eoc|> separators. This token-centric design supports both simple cross-modal pairs and complex interleaved datasets (MMC4-style zhu2023multimodal ), while preserving positional relationships and reducing redundant computing of shared media files. Core fields (“text”, “query”/“response”) also align with the popular formats of language foundation models, supplemented by bi-directional conversion tools for training ecosystems like LLaMA-Factory zheng2024llamafactory and ModelScope-SWIFT zhao2024swiftascalablelightweightinfrastructure (implementation details are in Appendix E.2).
Internal Adaptation with Industrial-Grade Optimization. The intermediate layer design enables us to adaptively take advantage of different underlying engines and systematically improve usability. We can optimize once and apply the enhancements with both execution modes. For example, our minhash-based Deduplicators achieve engine transparency and performance superiority, yielding 3.3x speedup over vanilla Ray with load-balanced union-find bts and hash-based aggregation.
Moreover, large-scale data processing faces schema/corruption errors that halt pipelines. Data-Juicer 2.0 supports sample-level fault tolerance through unified batched processing signatures and exception handlers that skip invalid batches (Fig. 7 in Appendix F.2), ensuring worst-case progress guarantees. Schema-compatible empty samples also maintain dataset integrity. To mitigate memory bottlenecks, we introduce streaming JSONL loading via patched Arrow backends, and automatic small file splitting aligning with Ray’s block distribution strategy. This reduces network overhead and improves CPU utilization. These optimizations are detailed in Appendix F.
4.2 Data-Juicer-Operators
Composable Processing Primitives. We extend Data-Juicer 1.0’s five atomic OPs (Formatter/Filter/Mapper/Deduplicator/Selector) with five compositional types: Grouper/Aggregator/FusedOP/ScriptOP/HumanOP. Using Strategy/Decorator patterns gamma1995design , they enable flexible algorithm encapsulation and runtime behavior extension (OP taxonomy in Appendix E.3.1). FusedOP optimizes batch-wise OP fusion (Appendix E.3.2), while ScriptOP integrates custom Python logic. The Grouper takes a Data-Juicer-Dataset as input and groups data samples into batches, which can then be input into the Aggregator for subsequent aggregation. All OPs follow a unified template method run() with automatic parallelism configuration, decoupling execution logic from runtime engines.
We design an abstract factory class that centralizes common functionalities (parameter preparation, serialization) while allowing for implementation of OP-specific execution logic through overridable methods (e.g., compute_stats() for Filters). This eliminates executor dependencies and enables easy standalone customization/inspection/testing of individual OPs as their functionalists are constrained to be implemented in a self-contained manner.
OP-wise Optimization. We incorporate several automatic adaptation features for Data-Juicer-Operator, aiming at balancing resource constraints and operational efficiency without requiring users to understand hardware specifics or implementation details.
We employ a dedicated Adapter class, which uses a probe_small_batch() method to systematically probe and analyze essential information by applying individual OPs on randomly sampled data in runtime. As a result, we enhance Data-Juicer 1.0 ’s greedy OP fusion with adaptive reordering based on estimated OP speeds. Faster OPs precede slower ones within commutativity constraints, optimizing end-to-end latency (validation in Appendix H.2.1). Moreover, using a uniform parallelism granularity across all OPs in a data pipeline can cause OOM issues for some and resource underutilization for others. In Data-Juicer 2.0, with auto-configuration, model-based OPs use GPU/quantization (e.g., vLLM), while I/O-bound OPs use hierarchical parallelism across batched processing, multiprocessing and multithreading, taking the concurrent opportunities between I/O and computation latencies. The implementation details for OP Adapter are in Appendix G.
OP Insight Mining. The combined effects of sequential OPs are not always additive, as validated in chen2024djsandbox . Existing tools like Data-Juicer 1.0 chen2024dj and Falcon refinedweb focus on coarse-grained metrics (e.g., data volume changes via Sankey diagrams) but lack fine-grained analysis.
To formulate better data recipes, Data-Juicer 2.0 tracks dataset statistics (e.g., perplexity) and semantic tags (e.g., image categories) after each OP execution. Built-in Analyzer leverages Filters and modality-specific tagging OPs to generate histograms of statistical distributions and semantic categories (an example is shown in Appendix, Fig. 8). Data-Juicer 2.0 automates metric comparison between consecutive OPs, producing reports that highlight significant lineage-level variations. For example, a sudden text-length reduction after applying a BLIP-2 li2023blip2 image-text matching Filter could indicate noisy captions requiring adjustment. These insights help users systematically evaluate OP impacts, optimize data recipes, and identify unintended correlations.
4.3 Processing Job Control
End-to-End Workflow Orchestration. Our Executor module integrates configurable data pipelines with monitoring/checkpointing capabilities, codified into reusable data recipes (the red box in Fig. 1). We also provide a sandbox suite that enables data-model co-development through template workflows connecting to model training/evaluation infrastructures chen2024djsandbox , allowing cost-effective exploration of data-compute effects before full deployment. More implementation details are in Appendix E.4.
Extensibility for Diverse Applications. Many effect-proven and illustrative workflows have been encapsulated in YAML recipes and maintained online dj-cookbook , catering to various vertical domains, such as multimodal data synthesis and persona-oriented dialog processing jiao2024img ; zhou2024humanvbench ; role_play_recipe . These facilitate interface exposure and reuse across different levels (Sec. 3), simplifying recipe routing and tailored processing solution generation by foundation-model-based agents.
5 Experiments & Insights
5.1 Experiment Setup
We evaluate Data-Juicer 2.0’s efficacy across three data scales: small (560K-2.24M samples), medium (5.6M-56M samples), and large (56M-70B samples), covering both multimodal and text-only processing. Our test suite executes five representative data operations per recipe across three compute engines (Standalone, Ray, MaxCompute) using Alibaba Cloud resources (1-100 nodes, 64-12,800 CPU cores). All worker nodes maintain identical hardware configurations for fair comparison. More implementation details are in Appendix H.1.
5.2 The Case of Small Scales
Performance Profile. As shown in Fig. 4d, when processing small-scale multimodal datasets, the standalone mode with Hugging Face Dataset is efficient and comparable to the Ray mode with a single node. Additional Ray nodes provide further but limited acceleration (speedup ratios between 138% and 226% with 4 nodes). For text-only datasets (Fig. 4a), the standalone mode remains efficient. However, for the Ray mode, 4x node increments yield smaller speedups (148%) or even increased processing times due to dominated I/O and communication costs. Thus, with Data-Juicer, processing datasets with hundreds of thousands of samples on a single machine is both efficient and cost-effective for most users.
Implications & Typical Application Scenarios. In this scale, Data-Juicer enables two critical capabilities for data-centric AI research: (1). Rapid recipe prototyping. It simplifies data-model co-design as exemplified by extensive sandbox experiments chen2024djsandbox , covering text-to-video generation, image-text pre-training, image-to-text generation for general image understanding, image captioning, and model prompt optimization. (2). Structured insight mining. It helps to support 5 open foundation model competitions on data filtering, augmentation, and synthetic data generation with 3,000+ teams dj-competition . Key lessons learned reveal that standardized and systematic actions provided by Data-Juicer (e.g., YAML recipes, probe sampling, visual analytics integration) accelerate data analysis and understanding compared to ad-hoc implementations.
5.3 The Case of Medium Scales
Performance Profile. When datasets are scaled to 56M samples, processing times increase significantly with the standalone mode, thus, its performance is omitted in Fig. 4e and Fig. 4b. Here, the Ray mode outperforms in all instances, demonstrating considerable speedups with increasing node counts, making it the recommended choice for medium-scale scenarios. Moreover, compared to native nodes of Elastic Computing Service (ECS, green lines), the Ray mode on Deep Learning Containers (DLC, red lines) is faster, saving 24.8% of processing time due to the Alibaba Cloud’s optimization dedicated to cluster networking.
Implications & Typical Application Scenarios. In this scale, Data-Juicer boosts data flywheels once we find high-quality data recipes. There have been many synthesis-based and ready-to-use data recipes built upon Data-Juicer like jiao2024img ; zhou2024humanvbench ; xu2025mindgym , where more compute investment brings larger-size datasets. Several lessons were learned from them: (1) Although the loss functions of foundation models are relatively standardized, we can flexibly inject preferred inductive bias with dedicated data synthesis, such as contrastive learning jiao2024img and data-difficulty based curriculum learning xu2025mindgym . (2) Foundation models emerge with expert-level knowledge, which can be used as proxy annotators, largely reducing manual labeling costs in benchmark construction zhou2024humanvbench .
5.4 The Case of Large Scales
Performance Profile. For datasets at the 70B-sample scale, professional cloud-scale distributed data processing products are advantageous. Users benefit from the vast cloud computing resources without dealing with the intricacies of setup and management. From our experiments, we recommend Ray-DLC for multimodal recipes and MaxCompute for text-only recipes at this scale. (1) For multimodal recipes, using 3200 Ray-DLC cores process datasets in 1780.86s and 7083.5s for 500x and 2500x dataset sizes, respectively, indicating good scalability. On the other hand, the MaxCompute engine requires 1.5 times more processing time using the same resources, due to the challenges of loading large-size multimodal data. (2) For text-only recipes (Fig. 4c), although the Ray mode benefits from additional cores even at the scale of ten thousand, MaxCompute is the fastest, requiring about 1/4 of the time and using 1/2 fewer cores. This is attributed to MaxCompute’s co-optimization of distributed computing and storage, a feature not as advanced in Ray’s implementation.
(a) Small-scale text-only data.
(b) Medium-scale text-only data.
(c) Large-scale text-only data.
(d) Small-scale multimodal data.
(e) Medium-scale multimodal data.
(f) Resource utilization comparison.
Figure 4: Processing time comparison across computing engines, dataset sizes, and workloads.
Impact of I/O and Data Splitting. I/O is critical for data processing workloads. For distributed processing, large-scale datasets need to be distributed to nodes, stored on shared disks like network attached storage (NAS), or on cloud storage services like Alibaba Cloud object storage service (OSS) and cloud parallel file storage (CPFS). All these introduce inevitable I/O and network communication costs. In our experiments shown in Fig. 4, we use Alibaba Cloud standard CPFS, compared to which NAS and OSS increase time costs by 20-30%. Furthermore, using another AI-oriented CPFS product with a 3x larger network bandwidth and optimized remote direct memory access to process the 12,500x dataset with 3200 cores in Ray mode takes 1,625s, 2.7x faster than the standard CPFS (4,396s). We then scale up the dataset to 125,000x, resulting in times of 14,617s and 7,611s with 3,200 and 6,400 cores, respectively, further demonstrating the efficiency and scalability of the system.
Additionally, the adaptive subset splitting capability for Ray mode, introduced in Sec. 4.1, offers 2$\sim$3x acceleration in our experiments. For example, with 100 Ray-DLC nodes and 12,800 cores, our subset splitting method reduces the processing time of the 12,500x dataset from over 5,000s to about 2,000s. Analysis of Ray’s resource utilization shows that its automatic block-splitting strategy incurs high network communication costs and low CPU utilization when many nodes handle a few dataset files. Our strategy pre-splits datasets according to size and node count, optimizing alignment with Ray’s features and enhancing efficiency. As seen in Fig. 4f, this strategy reduces peak network I/O from about 160MB/s to 60MB/s and achieves more consistent CPU utilization across all nodes.
Table 1: Deduplication time (minutes) with Ray nodes across varying dataset sizes and CPU counts (160 cores per node).
{tblr}
width = colspec = Q[140]Q[270]Q[240]Q[240], hline1,4 = -0.08em, hline2 = -, # CPU & 200GB Time 1TB Time 5TB Time
4*160 11.13 min 50.83 min 285.43 min
8*160 7.47 min 30.08 min 168.10 min
Large-scale Deduplication. We tested the MinHash-based RayDeduplicator with datasets from Redpajama together2023redpajama and Common Crawl chunks commoncrawl , sized at 200GB, 1TB, and 5TB, using CPU counts ranging from 640 to 1,280 cores. As shown in Table 1, RayDeduplicator efficiently scales with increased data size and computing resources, indicating its capability to handle large-scale deduplication tasks effectively. When the data size increases by 5x, the processing time increases by 4.02$\sim$5.62x. When the number of CPU cores doubles, the processing time decreases to 58.9$\sim$67.1of the original time. Notably, RayDeduplicator can process 5TB of data in 2.8 hours using 8 Ray nodes (8*160 CPU cores). As a comparison, NVIDIA’s NeMo Curator, which leverages cuDF cuDF and Dask dask , takes 1.8 hours to process 1.1TB of Redpajama data using 64 A100 GPUs (64*6,192 CUDA cores), as shown in nemo-dedup-results .
Implications & Typical Application Scenarios. The optimizations in Data-Juicer benefit many emerging paradigms to meet the needs of large-scale scenarios. For example, it powers enterprise foundation model training from Alibaba Tongyi and Alibaba Cloud’s production deployments, especially for TB-token pre-training and costly video/image processing for spatial intelligence. In addition, it helps to explore the data scaling law ge2024data and reinforced fine-tuning Trinity-RFT , where users need to efficiently process scalable feedback data for learning from experience with environment interactions. Moreover, the key lesson is that complex distributed system introduces non-linear scaling tradeoffs that depend on data modality and access patterns, and storage-compute-software co-design becomes critical beyond 10M samples.
5.5 More Ablation Studies on Runtime Adaptations
We conducted several experiments to evaluate the efficiency of enhancements from Data-Juicer core runtime. Overall, OP reordering along with OP fusion can save at most 70.22% processing time for some complex data recipes; Automatic GPU resource allocation helps to prevent OOM errors and saves at most 99% time when processing multimodal datasets with foundation models; Data processing with a batch size of 1000 reduces processing time by at most 84% in most scenarios; And OP-wise hierarchical parallelism speeds up the data processing consistently across all parallel levels (multiprocessing, batched processing, multithreading). One take-away message is that adaptive batching and resource allocation (e.g, via our hierarchical parallelism) can effectively mitigate hardware underutilization caused by irregular computation graphs of data pipelines for foundation models. More details about these ablation studies are presented in Appendix H.2.
6 Related Work
Data Infrastructure for Foundation Models. Modern data processing systems for foundation models need to reconcile conflicting demands between multimodal semantics and computational efficiency wang2023data ; bai2024survey . While existing frameworks excel at structured data analysis apache-spark or text-centric processing DolmaToolkit ; penedo2024datatrove , they struggle with three key gaps. First, current systems lack native support for cross-modal alignment - a critical requirement for quality-sensitive tasks like visual-language pretraining liu2023llava and multimodal synthesis podell2023sdxl . Second, traditional SQL-based optimizations become inadequate for pipelines dominated by neural network inference and fuzzy deduplication schuhmann2021laion , particularly when processing large unstructured records ikegwu2022big . Third, Python-centric workflows of foundation model practitioners Paszke-NeurIPS-2019-Pytorch conflict with JVM-based architectures of legacy big data systems.
Our Position. Next-generation data processing systems for foundation models require three key capabilities to bridge existing gaps: (1) Standard multimodal OPs that handle semantic relationships across vision, language, and audio modalities; (2) Hybrid execution strategies balancing local GPU acceleration with distributed processing for neural data transformations; (3) Python-native interfaces that abstract engine-specific complexities while maintaining compatibility with major computational backends (Hugging Face, Ray, Spark). As detailed in Appendix A, the proposed architecture addresses the tension between semantic-aware processing and scalable execution - a fundamental challenge when handling terabyte-scale multimodal datasets with heterogeneous compute requirements. Our implementation demonstrates how careful system design can bridge the gap between academic prototyping needs and industrial-scale data pipelines.
Difference to Prior Works. As user demands and applications evolved rapidly, the complexity of Data-Juicer 1.0 increased correspondingly. The need for heterogeneous data processing for more modalities with enhanced usability calls for an updated version. Besides, it showed shortcomings in programmability, such as fragmented development workflows across diverse underlying data frame frameworks. Unlike Data-Juicer 1.0 and other prior modality-specific and task-specific toolkits for or with large models Jennings_NeMo-Curator_a_toolkit ; lhoest-etal-2021-datasets ; curator ; penedo2024datatrove , we introduce a new composable OP system that generalizes across data types and training tasks while maintaining scalability optimizations. This contrasts with conventional big data systems ikegwu2022big that prioritize relational algebra over neural network integration, by emphasizing computational efficiency alongside semantic preservation.
7 Conclusions, Limitations & Future Works
Data-Juicer 2.0 emerges as a versatile scaffold in the evolution of foundation models, providing efficient and adaptive data processing solutions for handling the diversity and scale of modern datasets. Our re-envisioned architecture leverages multi-layered adaptability to coordinate different modules, operators, and runtime environments. Extensive evaluations reveal its high performance across diverse cloud-scale workloads. By open-sourcing this system, we aim to foster a vibrant community of contributors and users, encouraging collaborative development and driving innovation to underpin the next generation of foundation models.
There are several limitations in Data-Juicer 2.0. From the perspective of the computing system, future work includes further enhancing processing adaptability through model-driven agents gao2024agentscope , scalability by optimizing the single-node transmission bottleneck in Ray’s head node ray-head-worker , and efficiency by supporting GPU backend engines like NeMo Curator Jennings_NeMo-Curator_a_toolkit . From the perspective of societal impacts, we now introduce several operators for removing NSFW content, and most operators can only process data in English/Chinese. Expanding processing capabilities to cover multilingualism, data privacy, data safety, and so on, is one of the most important enhancement directions.
Appendix
[…]
1Table of Contents
Appendix A Discussions on System Challenges and Design Goals
A.1 Data Processing for Foundation Models
Foundation models, especially those beyond text, demands innovative data processing paradigms to address the complexity and diversity of large-scale training datasets [54, 79, 74, 5]. Despite existing systems for big data or text-only models [78, 21, 58, 32], there is a notable lack of open-source frameworks customized for foundation models due to unique design and implementation challenges:
Functionality Challenges from Data Types and Semantics. Data processing frameworks for foundation models primarily target semi-structured textual and multimodal datasets (e.g., JSONL, Parquet, MP4), often disregarding the strongly structured data types prominent in previous big data systems (e.g., database tables) [36, 32]. Significantly, the performance of foundation models depends on the quality, diversity, utility, and safety of data, necessitating a profound semantic understanding and cross-modal alignment to transform raw data into high-quality inputs, thereby enhancing model efficacy [40, 38, 70].
Efficiency Challenges from Data Operations and Pipelines. Conventional engines support a broad range of SQL queries for complex data operations [78, 60]. However, in the context of foundation models, the requirement for traditional grouping and aggregation operations is relatively minimal [5]. Instead, emphasis is placed on simpler operations such as Mapper and Filter, which rely heavily on deep learning models via local inference [57, 11, 73] or API calls [40, 43, 41]. Furthermore, individual records in these datasets can be exceedingly large, surpassing the capabilities of traditional big data engines [47, 31]. When combining multiple data operations into pipelines, it is crucial to maintain efficient data and transformation flows throughout the data lifecycle to prevent bottlenecks stemming from computational overhead, latency, and data integrity.
Usage Challenges from User Interfaces. Traditional big data systems typically use programming interfaces based on Java and Scala, supplemented by SQL. However, practitioners focusing on foundation models predominantly operate within the Python ecosystem, supporting both upstream and downstream repositories and workflows [49, 1, 55]. This shift necessitates designing user interfaces that specifically cater to the new programming needs of foundation model practitioners.
A.2 System Design Goals in Data-Juicer 2.0
Built upon Data-Juicer 1.0, we aim to address challenges identified in Appendix A.1 with the focus on the following new objectives:
Comprehensive Multimodal Processing Abilities. The first basic goal is to support extensive multimodal processing capabilities across diverse domains, such as AI-generated content, spatial intelligence, education, and entertainment. This involves meticulous processes such as collection, cleansing, annotation, and synthesis, which integrate perceptual and cognitive information across modalities such as video, image, text, and audio [10, 52, 2, 22].
Efficient Data Processing Operators and Data Flow. Comprehensive semantic operations are expected through advanced procedures that may be high-cost, such as deep learning model invocations, fuzzy deduplication, and clustering. For basic, high-frequency operations, we need to employ dedicated code optimization and speedup techniques like GPU computation or Cython [6]. Additionally, optimizing the loading, parsing, and storage of multimodal data, while minimizing overhead associated with linking, read/write operations, and network transmissions, is crucial for streamlining processing and enhancing efficiency.
Easy-to-Use Interface. To tackle usage challenges, Data-Juicer 2.0 pursues an intuitive, engine-agnostic abstraction of computational primitives, supporting Python, layered APIs, graphical user interfaces, and even natural language interactions. This design shields users from the inherent complexities of various computational engine implementations and manages the cognitive and usage costs to remain consistent, even as functionalities expand.
Adaptive and Scalable Processing. Last but not least, Data-Juicer 2.0 is designed to dynamically adapt to diverse computational environments and workloads, optimizing performance across frameworks including Hugging Face Dataset, Ray Dataset, and Alibaba Cloud MaxCompute, while considering their respective use cases. For example, OPs based on Hugging Face Dataset align well with large model community compatibility but require enhancements for scalable and distributed processing. Ray-based OPs offer excellent Python compatibility with low interaction overhead, yet they may face single-node bottlenecks for large-scale clusters involving data shuffles, such as those required by join operation. OPs based on Spark and MaxCompute necessitate careful resource management and schema considerations due to serialization between Python and the backend engines (e.g., JVM), rendering them less suitable for high-frequency computations.
In the subsequent sections, we will explore the design specifics and implementation of Data-Juicer 2.0 to illustrate how these objectives are achieved.
These design goals guide the development of robust operator functionalities and system implementation, tailored to meet the specific requirements of foundation model data processing. In the subsequent sections, we will explore the design specifics and implementation of Data-Juicer 2.0 to illustrate how these objectives are achieved.
Appendix B List of New Operators in Data-Juicer 2.0
The full list of new OPs in Data-Juicer 2.0 is shown in Table LABEL:tab:new_ops.
{longtblr}
[ caption = List of new OPs in Data-Juicer 2.0. The full OP list can be found in our document in GitHub repository., label = tab:new_ops, entry = none, ] width = colspec = Q[30]Q[200]Q[400]Q[85]Q[80], cell21 = r=12, cell141 = r=12, cell261 = r=27, cell531 = r=8, cell611 = r=5, cell661 = r=25, Modality Type OP Name (OP Type as the last suffix) OP Description Function Type Implementation Type
Image-only image_blur_mapper Blur images Augmentation Rule-based
image_remove_background_mapper Remove background of images Augmentation Model-based
image_face_blur_mapper Blur faces detected in images Privacy Model-based
image_aesthetics_filter Keeps samples containing images whose aesthetics scores are within the specified range Cleaning Model-based
image_aspect_ratio_filter Keeps samples containing images with aspect ratios within the specified range Cleaning Rule-based
image_face_ratio_filter Keeps samples containing images with face area ratios within the specified range Cleaning Model-based
image_nsfw_filter Keeps samples containing images with NSFW scores below the threshold Cleaning Model-based
image_shape_filter Keeps samples containing images with widths and heights within the specified range Cleaning Rule-based
image_size_filter Keeps samples containing images whose size in bytes are within the specified range Cleaning Rule-based
image_watermark_filter Keeps samples containing images with predicted watermark probabilities below the threshold Cleaning Model-based
image_deduplicator Deduplicates samples at document-level using exact matching of images between documents Deduplication Model-based
ray_image_deduplicator Deduplicates samples at document-level using exact matching of images between documents on ray Deduplication Model-based
Image-Text image_captioning_from_gpt4v_mapper Generate texts based on GPT-4V for images Synthesis Model-based
image_captioning_mapper Generate texts based on image-to-text models for images Synthesis Model-based
image_diffusion_mapper Generate images based on text-to-image models for texts Synthesis Model-based
mllm_mapper Use multimodal large language models for visual question answering tasks Wrapper of external tool  Model-based
sdxl_prompt2prompt_mapper Use the generative model SDXL and image editing technique Prompt-to-Prompt to generate pairs of similar images Synthesis Model-based
image_segment_mapper Perform segment-anything on images and return the bounding box values Wrapper of external tool  Model-based
sentence_augmentation_mapper Augment sentences using large language models Augmentation Model-based
image_text_matching_filter Keeps samples with image-text classification matching score within the specified range based on a BLIP model Cleaning Model-based
image_text_similarity_filter Keeps samples with image-text feature cosine similarity within the specified range based on a CLIP model Cleaning Model-based
phrase_grounding_recall_filter Keeps samples whose locating recalls of phrases extracted from text in the images are within a specified range Cleaning Model-based
image_pair_similarity_filter Keeps image pairs whose cosine similarity of features is within a specified range, based on a CLIP model. Cleaning Model-based
text_pair_similarity_filter Keeps text pairs whose cosine similarity of features is within a specified range, based on a CLIP model Cleaning Model-based
Video-only video_face_blur_mapper Blur faces detected in videos Privacy Model-based
video_ffmpeg_wrapped_mapper Wrapper to run a FFmpeg video filter Wrapper Tool-based
video_remove_watermark_mapper Remove watermarks in videos Cleaning Model-based
video_resize_aspect_ratio_mapper Resize video aspect ratio to a specified range Augmentation Rule-based
video_resize_resolution_mapper Map videos to ones with given resolution range Augmentation Rule-based
video_split_by_duration_mapper Split video by duration Augmentation Rule-based
video_spit_by_key_frame_mapper Split video by key frames Augmentation Rule-based
video_split_by_scene_mapper Split videos into scene clips Augmentation Model-based
video_human_tracks_extraction_mapper Extracts human tracks by linking face and body bounding boxes across frames Synthesis Model-based
video_human_demographics_mapper determines demographic attributes (gender, age, race) for face tracks by aggregating frame-level detections. Synthesis Model-based
video_human_description_mapper generates individual-focused videos from body bounding box tracks and processes it for appearance description and simple actions. Synthesis Model-based
video_facial_description_mapper generates face-focused videos from face bounding box tracks and processes them for emotion and facial description. Synthesis Model-based
video_active_speaker_detection_mapper detects active speaking by analyzing face track sequences alongside corresponding audio Synthesis Model-based
video_ASR_mapper generate the automatic speech recognition result for video Synthesis Model-based
video_speech_emotion_recognition_mapper Detect the emotion category of the speech in the video Synthesis Model-based
video_voice_demographics_mapper analyzes voice demographics (such as gender, age) in video Synthesis Model-based
video_aesthetics_filter Keeps samples whose specified frames have aesthetics scores within the specified range Cleaning Model-based
video_aspect_ratio_filter Keeps samples containing videos with aspect ratios within the specified range Cleaning Rule-based
video_duration_filter Keep data samples whose videos’ durations are within a specified range Cleaning Rule-based
video_face_ratio_filter Keep samples whose frame ratio containing faces is greater than a certain threshold are retained. Cleaning Model-based
video_motion_score_filter Keep samples with video motion scores within a specific range Cleaning Rule-based
video_nsfw_filter Keeps samples containing videos with NSFW scores below the threshold Cleaning Model-based
video_ocr_area_ratio_filter Keep data samples whose detected text area ratios for specified frames in the video are within a specified range Cleaning Model-based
video_resolution_filter Keeps samples containing videos with horizontal and vertical resolutions within the specified range Cleaning Rule-based
video_watermark_filter Keeps samples containing videos with predicted watermark probabilities below the threshold Cleaning Model-based
video_deduplicator Deduplicates samples at document-level using exact matching of videos between documents Deduplication Model-based
ray_video_deduplicator Deduplicates samples at document-level using exact matching of videos between documents on ray Deduplication Model-based
Video-Text video_captioning_from_audio_mapper Generate texts for videos according to their audio streams based on audio LLMs Synthesis Model-based
video_captioning_from_frames_mapper Generate texts for videos according to sampled frame images based on image-to-text models Synthesis Model-based
video_captioning_from_video_mapper Generate texts for videos based on video-to-text models Synthesis Model-based
video_tagging_from_audio_mapper Generate tags from audio streams extracted from the video. Synthesis Model-based
video_tagging_from_frames_mapper Generate video tags from frames extracted from the video. Synthesis Model-based
video_captioning_from_summarizer_mapper Generate texts by summarizing several types of generated texts (from video/audio/frames, tags from audio/frames, …) Synthesis Model-based
video_frames_text_similarity_filter Keep data samples whose similarities between sampled video frame images and text are within a specific range Cleaning Model-based
video_tagging_from_frames_filter Keep samples containing videos with given tags Cleaning Model-based
Audio-only audio_ffmpeg_wrapped_mapper Wrapper to run a FFmpeg audio filter Wrapper Tool-based
audio_add_gaussian_noise_mapper Add gaussian noise to audio. Augmentation Tool-based
audio_duration_filter Keep data samples whose audios’ durations are within a specified range Cleaning Rule-based
audio_nmf_snr_filter Keep data samples whose audios’ Signal-to-Noise Ratios are within a specified range Cleaning Rule-based
audio_size_filter Keep data samples whose audios’ sizes are within a specified range Cleaning Rule-based
Text-only calibrate_qa_mapper Calibrate question-answer pairs based on reference text Augmentation Model-based
calibrate_query_mapper Calibrate query in question-answer pairs based on reference text Augmentation Model-based
calibrate_response_mapper Calibrate response in question-answer pairs based on reference text Augmentation Model-based
extract_entity_attribute_mapper Extract attributes for given entities from the text. Synthesis Model-based
extract_entity_relation_mapper Extract entities and relations in the text for knowledge graph. Synthesis Model-based
extract_event_mapper Extract events and relevant characters in the text. Synthesis Model-based
extract_keyword_mapper Generate keywords for the text. Synthesis Model-based
extract_nickname_mapper Extract nickname relationship in the text. Synthesis Model-based
extract_support_text_mapper Extract support sub text for a summary. Synthesis Model-based
extract_tables_from_html_mapper Extract tables from HTML content. Cleaning Rule-based
generate_qa_from_examples_mapper Generate question and answer pairs based on examples. Synthesis Model-based
generate_qa_from_text_mapper Generate question and answer pairs from text. Synthesis Model-based
optimize_qa_mapper Optimize both the query and response in question-answering samples. Augmentation Model-based
optimize_query_mapper Optimize the query in question-answering samples. Augmentation Model-based
optimize_response_mapper Optimize the response in question-answering samples. Augmentation Model-based
pair_preference_mapper Construct paired preference samples. Augmentation Model-based
text_chunk_mapper Split input text to chunks. Synthesis Model-based
naive_grouper Group all samples to one batched sample. Synthesis Rule-based
key_value_grouper Group samples to batched samples according values in given keys. Synthesis Rule-based
entity_attribute_aggregator Return conclusion of the given entity’s attribute from some docs. Synthesis Model-based
most_relavant_entities_aggregator Return most relevant entities with the given entity from some docs. Synthesis Model-based
nested_aggregator Considering the limitation of input length, nested aggregate contents for each given number of samples. Synthesis Model-based
llm_quality_score_filter Keep sample with high quality score estimated by LLM. Cleaning Model-based
llm_difficulty_score_filter Keep sample with high difficulty score estimated by LLM. Cleaning Model-based
domain_diversity_selector Select samples based on the data’s domain diversity. Cleaning Model-based
Appendix C Showcase of Typical Multimodal Operators
The new OPs include some unique contributions to Data-Juicer 2.0, and others partly inspired by SOTA data processing methodologies for foundation models [57, 83, 26]. Below are three representative examples that showcase the diverse computational operations and requirements of these OPs.
phrase_grounding_recall_filter: This OP, newly developed by Data-Juicer 2.0, assesses alignment and consistency between images and textual descriptions. As illustrated in Fig. 5a, it identifies noun phrases in the text that refer to key entities. Subsequently, an open-vocabulary object detection model, such as Owl-ViT [44], attempts to detect corresponding entities within the image. The OP then calculates the recall of detected phrases to evaluate consistency, where a higher recall indicates better image-text coherence.
video_motion_score_filter: This OP quantifies video dynamics. As shown in Fig. 5b, it samples multiple frames at a specified frames-per-second (FPS) rate and computes optical flows. The average magnitude of these flows determines the motion score, with higher scores indicating more dynamic content. To accommodate various computational resources, both a GPU-based RAFT version [68] and a CPU-based OpenCV version [7] are available.
video_captioning_from_summarizer_mapper: This OP generates new video captions by combining tagging and captioning capabilities from various OPs. As illustrated in Fig. 5c, it uses six OPs from different angles: two for tagging and captioning the audio stream, two for static visual frame analysis, and two for dynamic video stream and information integration. A summarizer finally incorporates the three captions and two tag sets into a new caption. This composition yields more accurate and comprehensive captions by considering multi-dimensional and multi-perspective content.
(a) Example of phrase_grounding_recall_filter.
(b) Example of video_motion_score_filter.
(c) Example of video_captioning_from_summarizer_mapper.
Figure 5: Illustrative examples of new operators.
Appendix D User Interfaces
D.1 Low-level APIs
The framework’s core capabilities are exposed through Python-based programmatic interfaces, providing object-oriented logical encapsulation for both the fundamental Data-Juicer-Dataset, Data-Juicer-Operators and other runtime modules. This design delivers developers maximum flexibility and customizability. Processing workflows can be automatically chained by passing a series of OP instances to a loaded Data-Juicer-Dataset object (e.g., data.process([op1, op2])), enabling various operations to be performed on the dataset in a single pass. Additionally, the framework supports applying an instantiated Data-Juicer-Operators to a target dataset (e.g., op.run(data)), enhancing the reusability of OP instances. More details about the Data-Juicer-Dataset and Data-Juicer-Operators are introduced in Sec. 4 and Appendix E. This dual approach—chained processing and individual OP application—optimizes both efficiency and modularity in data processing tasks, while leveraging the inherent advantages of Python’s ecosystem.
D.2 RESTful APIs
Utilizing standard Python type hints, we provide one-click generation of high-performance web APIs, capable of automatically discovering, registering, and adapting OP classes and related tools. Users can rapidly initiate a web server supporting the Asynchronous Server Gateway Interface by simply executing a service script, eliminating the need for manual code writing. The asynchronous concurrency mechanism enables options such as lightweight background tasks and mitigates potential bottlenecks for endpoints that may experience prolonged network I/O blocking. Each OP is accessible via POST requests, typically executing the OP’s run() method as the endpoint. The target dataset path is passed through query parameters, with additional configurable OP parameters transmitted via JSON payload. Upon completion, the path to the processed dataset is returned. This invocation through Web APIs allows for a centralized host with distributed access, reducing deployment complexity. It also facilitates the separation of application logic from execution, potentially fostering the development and release of more applications built upon Data-Juicer. Importantly, the extensive customization parameters available in the programming API can be seamlessly passed through the Web API, maintaining full functionality without compromising usability, and facilitating serviceful calling by higher-level interfaces as introduced later.
D.3 Web & CMD tools
Figure 6: An example of config file for data recipe.
Thanks to the low-level and RESTful APIs, we further establish service-level capabilities across various scenarios in Data-Juicer 2.0, encapsulating scenario-specific processing solutions into built-in data recipes, described as YAML configuration files for end-to-end workflows (an example for video-data-synthesis is shown in Fig. 6). Centered around these data recipes, we provide user-friendly, highly encapsulated command-line interfaces as illustrated in Listing D.3, supporting automated lightweight installation and recipe-level invocation. These tools automatically gather and install the requisite of OPs within the recipe, perform automated OP fusion, and maximize adaptive utilization of computing resources while minimizing user cognitive load and operational cost.
{listing}
[!ht] {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]bash dj-install –config recipe.yaml # install dependencies only for this recipe dj-analyze –auto –dataset_path xx.jsonl # default OPs, or via specific recipe dj-process –config recipe.yaml # for both standalone or distributed Representative tools available in Data-Juicer 2.0.
Utilizing the aforementioned APIs and tools, Data-Juicer’s capabilities are seamlessly integrated into Alibaba Cloud’s visual modeling product, PAI-Designer [18]. As shown in Fig. 1, Data-Juicer-Operators are encapsulated as visual components within Designer, presented in a directory tree structure for easy access. Users can manage data by intuitive drag-and-drop manners, and import built-in or customized recipes in YAML format using graphical canvas interface. With Designer as entry point, users can activate MaxCompute cluster nodes [17] or containerized deep compute clusters (PAI-DLC) [16] directly for efficient cloud-native distributed data processing, leveraging Alibaba Cloud’s data storage solutions (e.g., Object Storage OSS [20], Cloud Parallel File System CPFS [15], and NAS [19]) along with automated performance monitoring and optimization.
D.4 Natural Language Interaction
User: Filter out too short text samples. The dataset path is input_data.jsonl Assistant: <thought>Filter out text samples that are too short based on the provided dataset path.</thought>
<function>run_text_length_filter</function>
<dataset_path>input_data.jsonl</dataset_path>
[…]
[ARGUMENTS] dataset_path: input_data.jsonl
[STATUS] SUCCESS
[RESULT] Filtered dataset path: output.jsonl
Assistant: <thought>The filtering has been successfully executed, and the filtered dataset is now at the specified path.</thought>
<function>finish</function>
<response>Your request to filter out too short text samples has been completed. The processed dataset is available at output.jsonl.</response>
Assistant: Your request to filter out too short text samples has been completed. The processed dataset is available at output.jsonl.
Querying and understanding over 100 diverse OPs have become challenging. Thankfully, advances in foundation models have transformed human-computer interactions. Research shows the synergy of reasoning and acting (ReAct) [77] in language models enhances task-solving. Using RESTful API services, we developed model-driven agents based on ReAct, enabling intuitive data processing through natural language. We adopted AgentScope [28], a multi-agent platform, for low-code integration, using prompt-optimized ReAct agents to align OP functionalities with our RESTful APIs. Built-in function preprocessing, response parsing, and fault-tolerance streamline development. Data-Juicer 2.0 thus allows flexible instruction input, interpreting user intents to execute appropriate data processing functions.
An example is shown above for interactive text length filtering using this agent, where users describe tasks with simple and vague language. Specifically, the user requests “filter out too short text samples from input_data.jsonl”. The agent’s ReAct prompt helps analyze and reason user intent, signaled by the <thought> tag. The <function> tag notes the chosen function, while <dataset_path> represents parameter mapping. Initially, a statistical analysis of text lengths is visualized, and the agent asks for filtering thresholds. After user input, filtering runs with parameters via RESTful API. Upon completion, the agent confirms the processed dataset path as output.jsonl, marks the task finished, and informs the user. Gray dialogs reveal the agent’s reasoning and actions, providing insights into the workflow. This example illustrates step-by-step user understanding, function selection, parameter population, and task execution, emphasizing transparency and automation of Data-Juicer 2.0 with agentic processing.
Appendix E Core Runtime Implementation Details
E.1 Execution Mode Configuration
The core runtime of Data-Juicer 2.0 is implemented in Python, featuring a flexible data processing pipeline characterized by two execution modes: a standalone mode for convenient single-node execution, based on the Hugging Face Dataset [36], and a distributed mode that offers scalability across multiple nodes, leveraging the Ray Data [45] and MaxFrame-DataFrame [17]. These frameworks offer diverse computational engines, each with distinct capabilities and suitable use cases, but they also come with heterogeneous programming interfaces and complex implementation details.
In Data-Juicer 2.0, we exploit the strengths of these diverse computational engines and their associated programming classes by abstracting a top-level Data-Juicer-Dataset class and devising a standardized data representation schema (detailed in the next subsection). The primary design principle employed is the Facade Pattern. The abstracted class provides comprehensive and unified development interfaces, facilitating seamless use across standalone, distributed, and cloud cluster environments, meanwhile optimized to support extensive multimodal data processing.
{listing}
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]python # Differernt input sources, such as from local files, dataset hubs from Hugging Face and Modelscope, and special types like arXiv, wiki, CommonCrawl, etc. data = DJDataset.load(src_file)
data = data.process(op1) # run a single operator data = data.process(op1).process(op2) # run multiple operators continuously data = data.process([op1, op2]) # run a list of operators
data.export(tgt_file) # tgt_file can be either local path or remote path Some basic interfaces of Data-Juicer-Dataset.
Moreover, following the principle of Template Method Pattern, the Data-Juicer-Dataset class establishes templated workflows for OP executing on datasets and provides several basic and unified interfaces, catering to users with diverse programming preferences, as demonstrated in Listing E.1. The complexities of underlying computational engines and cumbersome runtime issues related to instantiated dataset objects are abstracted away from these interfaces, ensuring transparency and ease of use for end-users and developers alike.
New functionalities can be added by merely implementing new OP classes without altering the internal logic of Data-Juicer-Dataset. The standalone and distributed execution modes automatically recognize and switch configurations based on runtime settings. Coupled with unified function signatures, the data processing can be consolidated into a batched processing mode, not only standardizing the data interface for OP processing but also facilitating the implementation of a robust sample-level fault tolerance mechanism. The internal adaptation features are elaborated in Appendix F.2.
E.2 Data Schema Implementation
{listing}
[!h] {minted} [ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ] json // >>> core contents: texts, dialogs, … // for general pretraining "text": "<__dj__image> desc of image 1 <|__dj__eoc|> desc of image 2 <__dj__image> <|__dj__eoc|>", // for post-tuning "query": "<query2>", "response": "<response2>", "history": [["<query1>", "<response1>"]], // <<< core contents
// $>$ extra data contents: multimodal data paths, … "images": [ "path/to/the/image1", "path/to/the/image2" ], // $<$ extra data contents
// $>$ meta infos and stats, which could be primitive or produced by Data-Juicer "meta": "src": "customized", "version": "0.1" , "stats": "lang": "en", "image_widths": [224, 336] , // $<$ meta infos and stats Illustration of data schema of Data-Juicer 2.0.
In the context of foundation models and post-tuning tasks, datasets are complex, originating from various sources and modalities. Target tasks are diverse, requiring different data formats and organizational structures. Integrating diverse data processing into a unified pipeline is challenging. Data-Juicer 2.0 adopts an extensible and intermediate data representation schema that intrinsically supports multimodal data types and flexible data field customization.
An example of a data sample is illustrated in Listing E.2. In the data schema, each dataset encapsulates column-stored samples represented by a non-recursive dictionary whose level-1 fields include three categories: (1) the core contents within customizable fields, including “text” that is usually used in the pretraining task, and “query”, “response”, “history” fields represent dialogs in post-tuning tasks, which are directly related to the pretraining or post-tuning procedures in the downstream usage of the dataset; (2) the extra data contents contain the path lists of multimodal data for multimodal datasets; (3) the “meta” information concerning this sample that stems from either its raw data or Data-Juicer’s Mappers for tagging, and the “stats” information calculated by Data-Juicer’s Filters (all Filter OPs consist of a compute_stats() method, utilized for subsequent analysis and conditional filtering).
For multimodal datasets, the content is primarily centered on the textual modality within the “text” field, capable of optionally including several text chunks split by customizable end-of-chunk tokens, which are <|__dj__eoc|> in default. Each chunk serves as a semantic unit, where the associated multimodal data within the same chunk relates to the same topic, thereby aligning with one another. Multimodal data (excluding text) are denoted by ordered customized special tokens in the text (e.g., <__dj__image>), storing entities at designated files, accessible through indexed file paths within the modality fields of samples to facilitate object sharing and elimination of redundant computing. For instance, in the example shown in Listing E.2, the first chunk in the “text” field split by the <__dj__eoc> token contains the first <__dj__image> token and a description to this image, which corresponds to the first image stored in the “images” field.
This token-centric representation elegantly aligns multimodal data while preserving positional information, friendly to current prevalent learning paradigms for foundation models employing next-token prediction tasks [5, 72]. And it’s compatible with both simple cross-modal pairing datasets (e.g., image-text pairs [14] and video-text pairs [75]), as well as complex interleaved multimodal datasets (e.g., MMC4 [83]), owing to its chunkable and special token design.
For post-tuning datasets, we provide several core fields in the core contents to represent the dialog datasets in our intermediate format, which is naturally aligned with Alpaca and Query-Response formats in the ModelScope-SWIFT repository [80] for smooth training of 400+ foundation models. As for other widely-used formats in well-known repositories like LLaMA-Factory [81], a suite of dataset conversion tools was developed to efficiently transform various popular multimodal datasets in them to and from the Data-Juicer schema, enabling users to process extensive existing models and datasets within Data-Juicer. These tools also serve as demonstrations for extending support to other uncovered datasets.
E.3 Operator Factory & Taxonomy
E.3.1 OP Design Principles
Operators are fundamental units responsible for executing processing functionalities, such as enhancing the accuracy of descriptive text or filtering out images containing Not-Safe-for-Work (NSFW) content. As illustrated by the middle orange box in Fig. 1, Data-Juicer defines five previous atomic OP classes: Formatter, Filter, Mapper, Deduplicator, and Selector; alongside four new types of compositional OPs: Grouper, Aggregator, FusedOP, and ScriptOP. The first five OP classes handle dataset format conversion, sample filtering, modification, deduplication, and selection, respectively. Following the Strategy Pattern, they are designed to encapsulate diverse algorithms that can be dynamically used to process the data. Each OP has a clearly defined role and can be interchanged or modified without impacting other system parts.
{listing}
[!ht] {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=, linenos, breaklines=true, bgcolor=lightgray!10, rulecolor=black, numbersep=5pt, ]python # Default approach processes the whole dataset sequentially data.process(op1).process(op2) # or data.process([op1, op2])
# Explict fine-grained processing at the batch level data.process(FusedOP([op1, op2], bs=1000)) # Internally for data_batch in data.next(batch_size): data_batch.process(op1).process(op2) Illustration of FusedOP in Data-Juicer 2.0.
Moreover, following the Decorator Pattern, compositional OPs are provided to enhance existing functionality without modifying the prevailing OP structure while dynamically adding data processing behavior to objects. The FusedOP enables explicitly grouping multiple atomic OPs to process data in a fine-grained manner within the same data batch, opposed to the default sequential processing across datasets as demonstrated by Listing E.3.1. The Grouper takes a Data-Juicer-Dataset as input and groups data samples into batches, which can then be input into the Aggregator for subsequent aggregation. For example, we can employ extract_entity_attribute_mapper and entity_attribute_aggregator following a key_value_grouper for meta-information extraction from textual input. Meanwhile, the ScriptOP allows users to incorporate existing Python files or execute code snippets, utilizing customized functionalities encoded within scripts like helper_func.py. This includes leveraging existing Data-Juicer command tools or integrating short Python scripts (e.g., lambda functions).
Together, these nine OP types provide robust expressive capabilities for end-to-end data processing solutions that can be embedded within a single YAML configuration file (as depicted in Fig. 6).
E.3.2 Unified Coordination of Logical Operations
In Data-Juicer 1.0, the logical operations of different OPs were coordinated within various executors (either standalone or distributed) rather than being bound to the OPs themselves. Disentangled from the executor’s scheduling interface, it becomes challenging to determine the execution logic of different OPs, making the development and extension of individual OPs less intuitive and lacking self-explanatory qualities. In Data-Juicer 2.0, several design principles are utilized to address this issue, including the Abstract Factory Pattern, Template Method Pattern, and Single Responsibility Principle.
Specifically, a top-level OP factory class is abstracted above the aforementioned fundamental OP classes. In this class, functionalities common to all OPs are extracted, such as preprocessing of instantiated parameters, support for serialization, and configuration of OP-aware runtime parallelism. Besides, a unified run() method is implemented, maintaining a consistent interface for integration and API calls. Furthermore, parallelism in multi-processing and multi-GPU is automatically configured and decoupled from specific OPs, ensuring transparency for end users and developers, as introduced subsequently in Appendix G. Lower-level OP classes define their own templated execution logic behind the run invocation. Taking the Filter class as an example, its core logic first engages the compute_stats() method to obtain statistical information based on specific metrics and then invokes its process() method to determine sample filtering based on thresholds.
On the one hand, this simplifies users’ understanding of OP types. Users can instantiate any OP and invoke it with a unified parameter signature using op.run(), thereby reducing the learning curve. The templated execution flow of various OP types is self-contained within base classes, eliminating dependencies on external executors to oversee invocation logic. On the other hand, by templating execution logic within base classes, developers can readily modify, extend, or implement new OPs. For instance, a developer aiming to customize an existing Filter does not need to rewrite a new class entirely but can inherit from a related existing class and override specific methods such as compute_stats() or others as required.
Regarding the naming and implementation of specific leaf OP classes, we adhere to extracting functionalities that are not tightly coupled with the OPs into common utility classes wherever feasible. This approach enables each OP to focus on its specific modalities and functionalities, facilitating a reduction in code complexity and enhancing clarity in understanding individual OP classes. Compared to previous implementations, the revised OP classes demonstrate easier inspection, integration, and testing. Users can utilize these robust OPs and seamlessly integrate them into their own tools or systems flexibly, both in source code or exposed RESTful API.
E.4 Control Panel Implementation
With the fundamental Data-Juicer-Dataset and Data-Juicer-Operators primitives established, we offer a series of control panel modules (the red box in Fig. 1) to organically combine them and accomplish data processing tasks. Executor encapsulates a series of standardized execution processes tailored for different standalone and distributed engines, leveraging modules such as Config and Monitor to accomplish end-to-end system configuration, data loading, analysis, processing iteration, data checkpointing, and more. Template workflows manage the complete data development lifecycle, including feedback, data storage, logging, and performance and operational monitoring.
To avoid the substantial costs from trial and error for data and model development in foundation model scenarios, we further develop a Sandbox suite in Data-Juicer 2.0 for data-model co-development, serving as a specialized intermediate layer connecting data processing jobs to numerous open-source infrastructures of model training and evaluation. The suite offers template workflows that extend beyond dataset-only development by incorporating cost-effective model training and evaluation signals, and quantitatively studying interplays among data, model, and compute. Users can easily conduct small, quick, and comparative experiments to find insights and superior data recipes, which can then be scaled to larger models and datasets, thereby optimizing the return on investment in data-model co-development. Additional details and empirically validated support can be found in [13].
Appendix F Optimization Details on Data-Juicer-Dataset
F.1 Engine-agnostic Processing
In Data-Juicer 1.0, the data processing pipeline is implemented using two distinctive execution modes: a standalone mode tailored for single-node operations and a distributed mode designed for scalability over multiple nodes. These modes exploit different dataset classes, each with a unique set of functionalities and interfaces. The default standalone execution mode employs the Hugging Face Dataset class [36], which provides a rich array of encapsulated functions, such as dataset.map() and dataset.filter(). It’s also equipped with configurable batch processing capabilities essential for various computational needs. In contrast, the distributed execution mode leverages the Ray Dataset class [45], which scales effectively across multiple nodes.
Despite both dataset classes using a storage format based on Apache Arrow [3], they exhibit significant differences in the behaviors and interfaces exposed. For instance, the Ray Dataset delineates individual and batched sample processing using separate methods: map and map_batches. It also allows to specify GPU counts for optimized scheduling. Meanwhile, the Hugging Face Dataset excels in supporting a broad spectrum of data modalities, such as image and audio. Moreover, the Hugging Face Dataset is typically applied in read-heavy data processing scenarios, such as in-memory tokenization and tensor reshaping, which are crucial for training deep learning models. For processing tasks in the context of foundation models, especially those involving synthesis operations, write-heavy procedures warrant attention. Here, the Ray Dataset provides flexible data exportation techniques advantageous to such tasks.
As highlighted in Sec. 4.1, we introduce a top-level Data-Juicer-Dataset class in Data-Juicer 2.0, along with common functions to bridge the variety of interfaces and implementation specifics across diverse computational engines. Besides the support of Hugging Face Dataset and Ray Dataset, our Data-Juicer-Dataset is also seamlessly extendable to support distributed computing within Alibaba Cloud ecosystems, thanks to the compatibility between ours and MaxFrame-DataFrame classes with intermediate in-memory formats like Pandas [48] and NumPy [30] or external-memory formats like Parquet. Thus, we develop unified wrappers that extract the core execution functions of different Data-Juicer-Operators into User-Defined Functions for the MaxCompute engine. Compared to Spark, MaxCompute is compatible in both syntax and runtime on Alibaba Cloud nodes and can be considered a commercially optimized version. An internal empirical comparison shows that MaxCompute SQL achieves 50% better performance than native Spark SQL.
The Deduplicators within Data-Juicer 2.0. Fuzzy deduplication is complex, involving a mixture of operations such as map, filter, group by, aggregate, and join [9]. The support and performance of these operations vary across different engines, especially in large-scale scenarios. To demonstrate the engine-agnostic feature of Data-Juicer 2.0, we use our minhash_deduplicator as an example, which supports the aforementioned three different engines. Users only need to specify algorithm-specific parameters such as jaccard_threshold and num_permutations. These parameters remain consistent regardless of the engine used, while engine-specific details and optimizations are completely transparent to the user. For instance, we utilize Ray Actors to implement our Ray-based deduplicator, starting with a load-balanced, distributed union-find algorithm [34] and introducing a hash-based aggregation process to enhance memory utilization and efficiency. This method avoids fragmented unions caused by Ray’s native groupby operation, which is computationally expensive in typical LSH implementations with traditional big-data engines [46]. As a result, we achieve a 3.3x speedup over our vanilla Ray version.
F.2 Fault Tolerance
In practical processing scenarios, datasets often contain schema-incompatible or corrupted data elements, such as improperly formatted JSON objects or damaged images that cannot be read. This issue becomes increasingly important with large-scale datasets, as processing tasks may extend over several hours or even days. In Data-Juicer 1.0, corruption of a single sample would halt the entire processing task, resulting in a waste of computational resources and the loss of already processed samples.
To address this issue, Data-Juicer 2.0 introduces a sample-level fault tolerance mechanism designed to enhance processing reliability by providing a worst-case guarantee. A unified exception manager is implemented to automatically capture runtime errors with customizable handlers during the processing of each OP. By default, dataset processing operations are performed at the data-batch level for a general handler. As a result, the system can easily bypass problematic samples by skipping the affected batch (as demonstrated in Fig. 7), while actively tracking and reporting these cases for subsequent debugging and correction. This ensures a seamless user experience and minimizes retry costs in scenarios involving large-scale data processing. Users have the flexibility to either discard these samples or mark them for future reprocessing.
Figure 7: Illustration of fault tolerance for “bad” data.
Of particular note is the fact that implementing this mechanism in Data-Juicer 1.0 presented challenges due to the rigid schema consistency mandates imposed by the underlying Hugging Face and Ray dataset classes. The newly refactored batching interfaces facilitate the unified construction of compatible empty samples based on schemas accessed during exception handling, thereby enabling seamless integration with valid dataset entries.
F.3 Streaming I/O and Subset Splitting
Memory constraints frequently emerge as a bottleneck in data processing tasks associated with foundation models. Memory demands can be substantial, and its precise usage is difficult to predict beforehand. Influential factors include the actual storage demands of objects such as text, image, and audio within individual dataset samples, as well as the storage demands of auxiliary or newly generated objects at runtime. These can stem from varying OP specifications, model sizes, synthesis data volumes, intermediate variables, the number of concurrent processes, and specific computational engines.
To effectively adapt to a diverse range of data processing scenarios with varying data volumes and available computational resources, we introduce streaming loading and data pre-splitting capabilities in Data-Juicer 2.0. They collectively facilitate improved computational resource utilization, and provide potential for a flexible programming space of hybrid stream and batch processing.
Firstly, Data-Juicer 2.0 offers a streaming loading interface, addressing the current lack of native support in the Arrow framework underlying Hugging Face and Ray Datasets for streaming JSONL data. As many foundation model datasets use JSONL, we developed an in-house patch to alleviate Out-of-Memory (OOM) issues.
Secondly, we develop a user-friendly script that automatically pre-splits the original dataset based on two observations: (1) The size limit of the underlying Apache-Arrow block, and (2) The inherent automatic block-splitting strategy in Ray. When there are tens of thousands of nodes but with only a few dataset files, Ray would split the dataset files according to the available resources and distribute the blocks of the dataset to all nodes, which brings a huge network communication cost and decreases the CPU utilization of each node. Thus, we split the original dataset into smaller 128MB files in advance automatically according to the dataset size and the number of distributed nodes, trying to adapt the features of Arrow and Ray for better performance. This approach reduces location and reprocessing costs associated with fault tolerance and helps mitigate network exchange overheads, especially beneficial in contexts involving large-scale multimodal data, as well as in scenarios that require handling global objects of Ray Actor in distributed modes.
Appendix G Implementation of Adaptation for Data-Juicer-Operator
This section explores the internal adaptations developed for Data-Juicer-Operator, which are crucial for optimizing resource allocation and user experience without requiring users to understand hardware specifics or implementation details. We implement several automatic adaptation features for resource management, aiming at balancing resource constraints and operational efficiency within Data-Juicer 2.0. These strategies include automatic OP reordering at the recipe level (Appendix G.1), automatic parallelism at the OP level (Appendix G.2), and automatic data insight mining to assess the impact of each OP on data samples, considering both upstream and downstream OPs (Appendix G.3). These features are encapsulated in a dedicated Adapter class, which uses a probe_small_batch() method to systematically probe and analyze essential information by applying individual OPs on randomly sampled data in runtime, with default sample size as min(1000, remaining_data_size).
G.1 Workloads-aware OP Reordering
At the recipe level, we introduce a new probe-based OP reordering strategy. In Data-Juicer 1.0, an OP fusion optimization was proposed to eliminate redundant computations for the Filter OPs, which involved three key steps: detection and grouping of fusible OPs, OP fusion, and OP reordering. The reordering strategy aimed to position more time-consuming OPs at the end of the group to process fewer samples to save time, as some were filtered out by preceding OPs. It is assumed that the fused OPs are the most time-consuming, and only the fused OPs are moved to the end of each group.
However, the assumption is not always correct and the prior reordering strategy omit the unfused OPs, rendering it greedy and suboptimal, especially when applied to diverse datasets and data recipes characterized by varied data distribution and OP orchestration. In Data-Juicer 2.0, we advance the reordering strategy to an adaptive approach, which is workload-aware and can be applied automatically to unfused OPs as well.
Specifically, before processing the full dataset, Data-Juicer 2.0 utilizes the probe functionality of Adapter to obtain estimated processing speeds for individual OPs relevant to specific input datasets. When processing the entire dataset begins, OPs in each group (including the unfused ones) are reordered based on the probed speeds and the commutativity of the Filter OPs. For the fused OP, assuming that there are $n$ fusible OPs in it and their probed speeds are $v_{i}$ where $i\in\{1,2,\cdots,n\}$, the estimated speed of the fused OP can be calculated as:
$v_{fused}=\frac{N}{T_{total}}=\frac{N}{\sum_{i=1}^{n}\frac{N}{v_{i}}}=\frac{1}{\sum_{i=1}^{n}\frac{1}{v_{i}}},$ (1)
where $N$ is the total number of samples to be processed by the fused OP and $T_{total}$ is the total time cost of it. Then, faster OPs are prioritized, while slower ones are deferred to later stages. This probe-based approach identifies the globally optimal reordering solution for each OP group, outperforming the suboptimal strategies of the previous version, as empirically validated in Appendix H.2.1.
G.2 Automatic Operator-wise Parallelism
In data processing for foundation models, it is crucial to recognize that different OPs require vastly different computational resources. Model-based OPs often need several gigabytes of GPU memory, while simple rule-based OPs may only need minimal CPU processing. Therefore, using a uniform parallelism granularity across all OPs in a data pipeline can cause OOM issues for some and resource underutilization for others. To address this challenge, we introduce several automatic mechanisms for OP-wise parallelism.
  • Model-based OPs that integrate substantial models require considerable computational time, potentially spanning hundreds of hours on CPUs for large datasets. To mitigate this, Data-Juicer 2.0 expedites these OPs by automatically leveraging CUDA and GPUs when available. Given the diverse memory requirements of large models, we utilize the Adapter component to conduct quick VRAM benchmarking prior to full-size dataset processing. This information is systematically assigned to the mem_required parameter of the respective OPs. During the execution of these OPs on datasets, Data-Juicer 2.0 continuously assesses the available GPU memory of the execution environment to dynamically and adaptively determine the optimal parallelism strategy. To further alleviate GPU memory demand, we also integrate quantization libraries such as vLLM [35] to enable efficient inference of leveraged models.
  • For non-model-based OPs, attributes such as cpu_required and mem_required are crucial. In Data-Juicer 2.0, users can specify a global parallelism level, typically aligning with the available processor count. Meanwhile, we calculate the adaptive number of processors to finely optimize the entire processing pipeline, aiming to maximize resource utilization (90% by default) as much as possible. For this purpose, the Adapter is also employed dynamically and instrumental in determining precise cpu_required and mem_required values at runtime.
  • For general-purpose and I/O-intensive OPs, Data-Juicer 2.0 enables batched processing, using a robust default batch_size parameter guided by performance saturation trends shown in Appendix H.2.3. Batched processing reduces I/O overhead, boosts efficiency per processor, and enhances overall speed. Additionally, we introduce hierarchical parallelism for OPs involving multimodal data I/O, such as image_aspect_ratio_filter. These OPs utilize multi-process and GPU parallelism, along with multiple threads, to handle data batches more efficiently, taking the concurrent opportunities between I/O and computation latencies.
G.3 Insight Mining Example
Here we provide an example of OP-wise insight mining in Fig. 8.
Figure 8: Illustration of OP-wise insight mining.
Appendix H Additional Experiment Details and Results
H.1 Implementation Details of Scalability Experiments
We start with a base dataset comprising 560k image-text pairs (about totally 5.6 million textual tokens) from the pertaining dataset of LLaVA [40]. We expand this dataset by factors of {2, 4, 20, 100, 500, 2500, 12500, 125000}, resulting in nine datasets scaling to 70B data samples. These datasets are categorized into three scales: small (1x, 2x, 4x), medium (4x, 20x, 100x), and large (100x, 500x, 2500x, 12500x, 125000x). For different scales, we prepare both multimodal and text-only data processing recipes, each containing 5 OPs. We run the recipes on these datasets using various computing engines (Standalone, Ray, MaxCompute) with different Alibaba Cloud resources, including ECS instances, PAI-DLC, and MaxCompute, spanning 1 to 100 nodes and 64 to 12,800 CPU cores. To ensure a fair comparison, we use the same worker node configuration, such as CPU frequency, across different computing engines.
H.2 Ablation Study on Runtime Adaptations
H.2.1 Automatic workloads-aware OP reordering
We evaluate the performance improvements due to our probe-based OP reordering on two recipes with different numbers of OPs. The simple recipe contains 5 OPs, with 2 fusible OPs, while the complex recipe includes 13 OPs, with 5 fusible OPs. We run these recipes on the base dataset used before (560k image-text pairs), comparing processing times for all OPs, fusible OPs, and other OPs.
Fig. 9 illustrates the results. Generally, OP fusion is effective for accelerating data processing, and automatic OP reordering offers further improvements based on OP fusion. Notably, OP reordering is more effective in complex data recipes (the right sub-figure), saving more time, especially for fusible OPs (46.09% v.s. 70.22%).
Figure 9: Processing time for simple and complex recipes.
H.2.2 Automatic GPU resource allocation
Efficient model inference on GPUs is critical for the processing speeds of model-based OPs. We select four image-related OPs, including an image-only OP and three image-text cross-modality OPs, integrating various models that require different GPU memories. Using a test dataset of 10k image-text pairs, we process these OPs on a machine with 64 processors and 8 Nvidia A100 GPUs, and compare the times with and without GPU support in Data-Juicer 2.0. Table 2 shows these comparative results.
Table 2: Multimodal OPs processing time comparison between CPU cores and GPUs on 10k samples. “np” denotes the number of processors.
{tblr}
width = colspec = Q[520]Q[100]Q[40]Q[150]Q[150], cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=2, hline1,7 = -0.08em, hline3 = -, hline2 = 4-5, Multimodal OPs & VRAM np Processing Time
CPU GPU
image_captioning_mapper ~16GB 32 8668.87s 305.44 s
image_diffusion_mapper ~8GB 64 100 h ~1 h
image_text_similarity_filter ~1.5GB 64 73.04 s 35.84 s
image_nsfw_filter ~1GB 64 102.59 s 39.74 s
As the results show, using GPUs saves at least 50% of processing time for all selected model-based OPs, especially for large, slow models like BLIP-2 [37] used in image_captioning_mapper and SDXL model [56] in image_diffusion_mapper. Compared to the CPU version, GPU usage and adaptive resource allocation are much more efficient and necessary. Due to the GPU memory limit of around 80GB, we can only allocate at most 4 models of image_captioning_mapper OP on a single GPU, so the number of processors for this OP is automatically reduced to 32, preventing OOM errors.
H.2.3 Batched data processing
We compare the efficiency of single-sample processing in the previous Data-Juicer 1.0 and batched processing in Data-Juicer 2.0. We use the previous 560k multimodal dataset, processed by 16 processors. Each processor handles about 30k samples. We select four recipes representing different scenarios: Filter-Heavy, Mapper-Heavy, Text-Heavy, and Multimodal-Heavy. Each recipe contains four OPs of the specified “heavy” type and one of another type. We run these recipes with five different batch sizes and summarize the results in Fig. 10a. From these experiments, we conclude: (1) Batched processing is always more efficient. Larger batch sizes consistently speed up data processing in all scenarios, reducing processing time by 84%. (2) Efficiency gains from larger batch sizes plateau beyond a certain threshold. Specifically, batch sizes of 100 or more yield similar benefits. (3) A batch size of 1000 is recommended, generally showing the most efficient processing in our experiments. Consequently, 1000 is selected as the default batch size in Data-Juicer 2.0.
(a) Different batch sizes and recipes.
(b) Different parallelism levels.
Figure 10: Data processing efficiency of OP-wise hierarchical parallelism. “bs” denotes the batch size, and “nt” dennotes the number of threads.
H.2.4 Automatic OP-wise hierarchical parallelism
To verify the efficiency of the automatic OP-wise hierarchical parallel strategy, we select the multimodal OP image_aspect_ratio_filter as an example. We conduct experiments with various processor counts, batch sizes, and thread numbers, covering three parallel levels: multiprocessing, batched processing, and multithreading. This OP processes approximately 560k image-text pairs. Fig. 10b demonstrates the time consumption and efficiency gains.
As depicted, significant speed improvements are consistently achieved across all parallel levels. Beyond the multiprocessing and batched processing strategies examined earlier, multithreading further reduces processing time. This OP benefits from multithreading due to its intensive I/O procedures (reading images), which balance I/O and CPU utilization. Most other multimodal OPs also require heavy data read/write operations, thus benefiting from the multithreading strategy.